A Bibliography of Literate Programming

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

07 January 2024 Version 3.39

Abstract

This bibliography records books and articles on the subject of literate programming, a term coined by Donald Knuth in [Knu84].

Title word cross-reference

= [Sar06]. ^{2} [Sar06]. · [Sar06].

0.87b [Brixx].

1 [CDL95, Pre95].

2 [Kre89, Pre95, Sew87a]. 21st [Ano94]. 268 [Chi92a].

3 [Ruc15].

5th [EB04, NBC92, SMEN96, USE97].

6th [Lio96].

’88 [Gia89]. 8X [AO90].

90 [AAK01]. ’93 [Ost93b]. ’94 [BGG^{+}94, Pur95]. ’96 [IEE96]. ’97 [USE97].

Abertay [SMEN96]. Abstract [BG87]. Abstraction [SC93b, Shu93, SC93a]. Abstraction-Oriented [SC93b, SC93a]. Accept [Hüs06]. Accept-and-Reject [Hüs06]. ACM [ACM93, Ano94]. ACM-SIAM [ACM93]. acquisition [Sha04]. Ada [Sen92, WB89]. adapted [GS86, Ken88, Lev87a]. Adelaide [NBC92]. Advanced [RW96]. Agile [PKB04b, EB04]. Aid [Hur96]. ALDES [Kre89]. ALDES/SAC [Kre89]. ALDES/SAC-2 [Kre89]. algebra [Kre89, Pag07]. Algebraic [vDHK96, Gia89]. Algorithm [Han95, Hüs06]. Algorithms [Knu00, Knu93b, Ruc05, ACM93]. Analysis [CS97, Knu00, Lei02a, Lum98, Oma88]. annotated [SS91b, SS91a]. Announcing [Den87]. Annual [ACM93, Ano94, BN93, IEE90, IEE96, USE97]. AOPS [Shu93, SC93b, SC93a]. APL [Dic93, NSW93]. APLWEB [Dic93]. Application [Aki92, AAK01, BP93]. Applications [RW96, NBC92]. Applied [BC90d]. Approach [vDHK96, Don01a, Kim97a, Lev87a, SW96]. April [Sit96]. Architectures [Kim97a, Kim97b]. arithmetic [AAK01]. Arizona [Joy94]. Array [Mit88b]. Array- [Mit88b]. Art [Ruc15]. Aspects [VN02]. Assessment [Van90b]. Assignments [Hur96]. Assurance [IEE96, MP96]. Australasian [Ros96]. Australia [Ros96]. authoring [GS05]. Automatic [Bri96a, Bri96d, CM92]. Automatically [Ait96]. available [Bzy95]. Avenarius [Kro90]. Axiom [Pag07].

Back [Mac98]. Balanced [Pfa04]. Based [DDPA96, DPDD96, KC02, Mot90, Ram94a, MGCR96, NTW91, MG94]. basierend [Fra97]. Be [Ram92, Ram91]. beyond [Lei03b]. bibliography [Sab94, SS91b, SS91a]. Bigot [Syr99]. Binary [Pfa04]. Bonn [FG96]. Book [OC90a]. Boston [USE97]. Breakpoint [Ram94a]. Bulletin [BP93]. Business [SMEN96].

C [Thi89, FH95, GS86, Han97, Hol90, Hym90, Lev87b, Lev87a, Pla92, SP92, Thi84, Wit95, van95]. Card [Chi93]. CASCON [BGG^{+}94]. case [Hol03, Pep91, PKB04a, Aki92]. CASEtool [Car95]. Castle [SMEN96]. CDS [Zen91]. change [NSW93]. changefiles [AH86]. Changing [Vee17]. character [Sab94]. chemistry [QW05]. CIP [BW91]. CIP-Project [BW91]. Clark [Ano10c].

Cleanroom [AMS97, Nor93]. clinical [GB06]. Cnest [KC92]. Co [SD95]. Code [Goo07, Hen97, Lio96, WB89, Thi03]. Coders [Sei09]. Coding [Vee17, Sab94]. Combinatorial [Knu93c, Knu93b, MN99]. Comment [Big99, Ait96]. Commentary [Lio96]. Common [VHG87, BKM86]. Communicating [Don01a, Don01b]. COMPASS [IEE96]. Compendium [Lun11]. Compiler [FH95, Hol90, Pre95]. Complex [Ram92, Ram91]. Complexity [SS92]. compliance [Sen92]. Comprehension [IEE94, JAC94]. compression [Sab94]. computation [Gia89]. Computational [NBC92, Sab94, LLAK05]. Computations [Lei03b]. Computer [BM86, IEE96, Joy94, Knu03a, Ros96, Ruc15, Vee17, Kre89, NTW91, Pag07, Knu86c]. Computing [Knu93c, SDDD12, CM92, MN99]. Concept [Lun11]. Concepts [DS97, MM94]. Conference [IEE90, IEE96, Mau95, NBC92, Pur95, Ros96, SMEN96, Sit96, EB04, FG96, Mah96, Ost93b]. Constraint [Hüs06]. containing [Ram96]. contemporary [PKB04a]. Control [Bet05]. Conventional [BC90d]. conversion [Sab94]. Correctness [Ram94a]. Cosmetic [OC90b]. Courses [Wit96a, Wit96b, CDL95, Wit96c]. craft [Goo07, Sei09]. Creating [Han97, Rou05, BP93]. CS [CDL95]. CS/1 [CDL95]. Cscope [KC92]. CuPit [Pre95]. CWEB [KL93, Wit95, Lev93, Thi86]. CWEBx [van95]. cycle [BC90b].

D [Bir92]. Data [Lei02a, BG87]. Debugging [BDM97]. Default [pt98]. Definition [Wai93]. Demonstrating [Sen92]. derivation [Pep91]. description [AAK01]. Design [BM86, BA86, BC90d, BC90e, FH95, Hol90, Ska98a]. Designing [Roy06]. Developers [DN07]. Development [Ost93a, Car95, Kre89, LLAK05]. developments [SW96]. dialogue [PG02]. diff [Thi89]. Difference [Thi89, VL89]. Digital [Knu99]. Discrete [ACM93, Knu03b]. display [Sab94]. Division [Han94, Han95]. Do [Kim97b]. doc [Ait96]. Docbook [Ano10b]. docstrip [pt98, Lun11, MDB92]. Document [Ano17, Kuh06, Fil08, Pre95]. Documentation [Ait96, Ano10b, BA86, Knu83a, KL93, VN02, WB89, Bal89, CS96, Ska98a, VN05]. documenting [Bal89]. documents [Coa98a, Lei03b, Sla90]. Donald [Ruc15, Adv00, DS94, Thi89]. Download [Ska98c]. Drawing [Gur94, Sof97]. DSSSL [Kim97b]. DTD [Coa98a]. Dudhope [SMEN96]. Dundee [SMEN96]. Dynamic [Lei02a].

econometric [MR09]. ED [Mau95]. ED-MEDIA [Mau95]. Edition [Lio96]. Editor [Mot90, Rea02b, Big99, Syr99]. Education [Joy94, Pur95, Ros96]. Educational [Mau95]. Einsatz [Sch92]. Eitan [Sof97]. electromagnetics [LLAK05]. Eleventh [IEE96]. Elucidative [Nør99b, VN02, Nør00]. Emacs [Chi93, MC91]. encryption [Sab94]. End [DN07, Mac98, PG02]. End-User [DN07, PG02]. Engaging [MM94]. Engineering [AMS97, BC90b, EB04, Ham95, Ost93b]. English [Mil94]. Enhance [PKB04b]. Enhanced [BM86]. Enterprise [RW96]. Entwurf [Kuh89]. Environment [BC87, BG92, Bro88a, BC90c, DDPA96, DPDD96, MGCR96, Mot90, Nør99b, SDDD12, Aki92, Hou92, KC92, MG94, Nør00, PG02, RS89]. Environments [Mit88b, Ost93a]. Eqn [Ano10c]. equations [Ram96]. Errata [Chi92a, Knu91]. everyday [GB06]. evolution [VN05]. Example [CH95c, CH95b, Pap90, Fil08, Knu82, QW05]. excellent [Goo07]. Executable [Bal89, Wai93]. Expanding [BC87, Ham88, VHC88]. Experience [RB98]. Experiences [Thi86]. Experiment [Ost93a]. experiments [NTW91]. Explaining [Thi03, Roy06]. Expressions [Ham88, Ram98, VHC88]. Extensible [Mot90]. extension [AAK01]. extensions [Bec86]. Extreme [EB04].

F2X [Don01b, Don01a]. F95 [Don01b]. Factors [BM90]. fifth [Joy94]. Figlet [Ano10a]. File [Thi89, VL89]. First [Lin89a, Ros96]. Fixed [Knu82, Knu91]. Fixed-point [Knu82, Knu91]. Florida [Sit96]. FOREET [BA86]. formal [SBR94, SW96]. Formalization [And02]. format [BP93, Kuh06]. formatter [vAK92a, vAK92b]. FORTRAN [AAK01, AO90, BA86]. Fourth [Sit96, ACM93]. Framework [Ano17, Zuk97, MR09]. Frameworks [Ano98]. Free [Ano10a, Ano10b]. Freetype [Ano10a]. functional [GS05, PL92]. FunnelWeb [Coa98a, Wil92, Wil00]. funnelweb.dtd [Coa98a].

Gaithersburg [IEE96]. Garmisch [EB04]. Garmisch-Partenkirchen [EB04]. Generalized [Ham88, VHC88]. Generating [Ait96]. Generation [Lei02a]. geometric [MN99]. Germany [EB04, FG96]. GIPSE [PG02]. glue [Knu82, Knu91]. GNU [Chi93, MC91, Ano10b]. Good [SC94]. GraphBase [Knu93b, Knu93c]. Graphite [Ano10a]. Grenzen [SD95]. Groff [Ano10c, Ano10a]. Gurari [Sof97].

hashing [Sab94]. Hawaii [IEE90]. headers [pt98]. held [NBC92]. Human [BM90]. Hypermedia [Mau95]. Hyperstructure [Ost93a]. Hypertext [DDPA96, DPDD96, Ost93b, Ost95, SP92, BC90a, Car95, Cze90, Øst93c, Par96].

IDE [Rac12]. Idee [Fra97]. II [Don01b, Lei03a, Lei03c, Mor94]. Implementation [FH95, Mit88b]. Implementations [Han97, Ram94a]. Implementierung [Kuh89]. Implementing [PL92, Lun11, Zuk97]. Improved [OC90a]. improving [Dun95]. Increasing [MP96]. Independent [Rac12, Ram89, VR89, vAK92a, Coa98a, vAK92b]. Indexes [Knu94]. input [Sab94]. Institute [IEE96]. Instructions [RF97]. Integrated [BGG^{+}94]. Integrating [AMS97, DN07]. Interactive [Bro88a, BC90c, Sla90]. interaktiven [Fra97]. Interface [HLR97, Cze90]. Interfaces [Han97]. Internal [VN02]. International [Gia89, IEE90, NBC92, Sit96, EB04, FG96, Mah96]. Interscript [Ska98a, Ska98b]. Intersections [Sha04]. Interview [Adv00]. Introducing [RW96]. Introduction [Chi92b, CH95a, Kyr95, Lin89b, Pfa04]. Introductory [Wit96a, Wit96b, Wit96c]. ISSAC [Gia89]. Italy [Gia89]. Iterative [Don01b].

James [Ano10c]. Java [Ait96, Dwe00, Sah02]. javadoc [Ait96]. Joe [Ano10c]. Joint [FG96, Mah96]. Jr. [Rob13]. Julia [Pas17]. July [Gia89, NBC92, Ros96, SMEN96, USE97]. June [EB04, IEE96].

Knowledge [DN07]. Knuth [Bir92, Adv00, DS94, Ruc15, Thi86]. Konzeption [Fra97]. Konzeptionelle [Sch92]. Krommes [Kro90].

Language [Ano10c, BC90e, vDHK96, Ram89, SDDD12, VR89, Wai93, vAK92a, GS05, vAK92b]. Language-Independent [Ram89, VR89]. Languages [Ano94, Knu03a, PL92]. LEDA [MN99]. Lehrerausbildung [Fra97]. length [Han94]. Leo [Rea02a, Rea02b, Swa98].

Lesk [Ano10c]. Letters [Big99, Syr99]. Lex [SM96]. Library [Pla92, LLAK05]. License [Ano10b]. life [BC90b]. life-cycle [BC90b]. Lindsay [Thi89]. Linear [Don01b, Ram96]. Linger [Han95]. Lions’ [Lio96]. LIPED [BG92]. Lisp [Hou92, Ram88]. Lite [MGCR96]. Literacy [DN07, Vee17, Sha04]. Literate [AMS97, vA90, Ano98, Ano10b, AO90, BC87, BK86, BKM86, Bir92, BG92, Bri93, Bri96a, Bri96b, Bri96c, Bri96e, Bri96d, Brixx, Bro88a, BC90c, BC90d, BC90e, Bzy95, Chi92b, Chi92c, CS96, CS97, Coa98b, CH95a, CH95c, CH95b, CB91, Cov98, Den87, DDPA96, DPDD96, DS94, DS97, Dic93, DN07, Dug93, Dun95, Dwe00, Gur94, Ham95, Ham88, HS98, Hen97, Hur96, Hym90, Jac87, Joh96, JJ97, KC02, Kim97a, Kim97b, Kna96, Knu83b, Knu84, Knu92, Knu93a, Knu94, Kuh89, Kyr95, Lec85, Lee94, Lei02a, LH07, Lev87b, Lev93, Lin92, Lin89a, Lin89b, MM94, Mit88a, MGCR96, Mot90, Ost93a, Øst93c, Ost95, Pap90, Par96, Pep91, PKB04b, QW05, Ram91, RM91a, RM91b, Ram92, Ram94b, Rea02a, Rea02b, RL03, Rou05, RB98, Sah02, SP92, Sch92, SDDD12]. Literate [Sew89, SC93b, SC94, SBR94, Sof97, SM93, SM96, Thi84, Thi86, Tun89b, VHG87, VJW87, VHC88, VL89, VR89, Van90a, Wit95, Wit96a, Wit96b, Van90b, Zen91, van95, vAK92a, And02, AAK01, Bro88b, BC90a, BC90b, BP93, CM92, CDL95, Chi10, Fox90, Gur90, GW90, GW91, Hou92, KH91, KH92, KC92, LLAK05, MP96, MG94, PH09, PKB04a, Pre95, Ram88, RS89, Roy06, Sab94, Shu93, SC93a, SW96, SS91b, SS91a, Tun89a, Wit96c, Wu90, vAK92b, Chi92a, Fra97, SD95, Moo03]. Literate-Programming [CB91, Ram92, AAK01, LLAK05, Fra97]. Lizards [Swa98]. Logic [DDPA96, DPDD96, FG96, Mah96]. London [BN93]. Long [Han95]. Look [Lin89a]. Lout [Ano10a]. LyX [Gor08].

M [Sof97, Mil94]. MA [USE97]. Machine [RF97]. Macro [Ano10c]. Maintaining [Rou05]. Maintenance [OC90a]. make [NTW91]. Makeindex [Ano10c]. Making [RW96]. MANGLE [Sew87a]. Manipulation [Bri96b]. manner [CDL95]. Manual [Lei02c, MC91, Spi91, Wil92, Wil00, van95]. March [Joy94]. Marking [Hur96]. Maryland [IEE96]. MasPar [Pre95]. Masses [DN07]. Mathematician [Knu93a]. Mathematics [Knu03b, Ruc05]. Matrix [Bri96b, Bri96c]. Maxtal [Ska98c]. Measuring [SS92]. mechanism [Dun95]. MEDIA [Mau95]. Meeting [BN93]. meets [Moo03]. Method [Tun89b, Tun89a]. Methoden [SD95]. Methods [BW91, DS97, PKB04b]. Mike [Ano10c]. Mills [Han95]. Minefield [Han94]. Mini [Knu94]. Mini-Indexes [Knu94]. Mixing [Lei02b]. MMIX [Ruc15]. mode [Chi93, MC91]. Model [HS98]. Modelling [RB98]. Modern [MM94, Par96, Knu86c]. Modula [Sew87a]. Modula-2 [Kre89, Sew87a]. Möglichkeiten [SD95]. MP [Pre95]. MP-1 [Pre95]. MP-2 [Pre95]. MP3 [Ruc05]. Multi [SDDD12]. Multi-Language [SDDD12]. Multimedia [Mau95]. Multiple [AH86, Han94]. Multiple-length [Han94]. Multiplication [Bri96c].

National [IEE96]. Need [Ram92, Ram91]. nonlinear [Ram96]. Notation [Wit96a, Wit96b, Wit96c]. November [Ost93b, Pur95]. Noweb [Ano10c, Kyr95, Ano10a, Mac98, JJ97]. Nroff [Ano10c]. Numerical [Don01a, Don01b]. Nuweb [Bri93, Brixx].

Object [MM94, LLAK05]. Object-Oriented [MM94, LLAK05]. odfWeave [Kuh06]. OOP [Don01a]. open [Pag07, Kuh06]. Operators [Ram98, Ram96]. Oppermann [Kro90]. ordering [Sab94]. Oriented [MM94, SC93b, LLAK05, Shu93, SC93a]. Orlando [Sit96]. Ossanna [Ano10c]. Otago [Pur95]. Outlines [Rea02b]. output [Sab94].

Package [Kuh06, Lei03a, Lei03c, Pap90]. Packages [Rou05]. Pages [Nør99a]. paper [Big99]. Papers [Joy94, Knu00, Knu03a, Knu03b, BW91]. Paradigm [CB91, OC90a, OC88a]. Paradigms [Swa98]. Parallelization [Bri96a, Bri96d]. Part [CH95c, CH95b, Mor94, Don01a, Don01b, Lei02b, Lei03a, Lei03c]. Partenkirchen [EB04]. Pearls [BK86, BKM86, BG87]. Permutations [Hüs06]. Perspective [VN02, Ham95]. perspicuous [Hol03]. Peter [Syr99]. Phoenix [Joy94]. Pic [Ano10c]. Plain [Bzy95]. Platform [Knu93c, Rac12, Knu93b, MN99]. Platform-Independent [Rac12]. point [Knu82, Knu91]. Polish [BP93]. Possibilities [RW96]. Postfix [Ram98]. pp. [Chi92a]. practice [GB06, Goo07, RL03]. Practices [SC94]. Practitioner [Chi92a, Chi92c]. Prefix [Ram98]. preliminary [GW90, GW91]. Presentation [BM86, SBR94]. Principles [Ano94, BM86, AAK01]. Printing [VHG87]. problem [Dun95]. Proceedings [ACM93, Ano94, BGG^{+}94, BN93, IEE90, IEE94, Mau95, Ros96, SMEN96, Gia89, IEE96, Mah96, NBC92, Pur95, Sit96, EB04]. processes [EB04]. Processing [Jac87, VJW87, Sab94, Sla90]. Processor [SM96, Sew87b]. Program [BM86, BKM86, IEE94, Knu86a, Knu86b, Ost93a, Sew89, VL89, Van90a, VN02, BK86, JAC94, Lum98, MDB92, Pep91, VN05]. Programmierumgebung [Fra97]. Programming [AMS97, Ano94, Ano98, Ano10b, Ano17, AO90, BC87, BK86, BKM86, BG87, BG92, Bri93, Bri96a, Bri96b, Bri96c, Bri96e, Bri96d, Brixx, Bro88a, BC90c, BC90d, BC90e, Chi92a, Chi92b, Chi92c, Coa98b, CH95a, CH95c, CH95b, CB91, Cov98, Den87, DS97, Dic93, DN07, Dwe00, Gur94, Ham88, HS98, Hen97, Hur96, Jac87, JJ97, KC02, Kim97a, Kim97b, Kna96, Knu83b, Knu84, Knu92, Lec85, Lee94, LH07, Lev87b, Lev93, Lin89a, Lin89b, MM94, MGCR96, Mot90, Nør99a, Nør99b, OC88b, Oma88, Ost93a, Ost95, Pap90, PKB04b, RM91b, Ram92, Ram94b, Rea02a, Rou05, Ruc15, RB98, SP92, SD95, Sch92, SDDD12, Sew89, SC93b, SC94, ST90, SM93, SM96, Swa98, Thi84, Thi86, Tun89b, VHG87, VJW87, VHC88, VL89, VR89, Van90a, Vee17, Wit95]. Programming [Wit96a, Wit96b, Van90b, vA90, vAK92a, And02, AAK01, Bro88b, BC90a, BC90b, BW91, BP93, CM92, CS96, Chi10, Dug93, Dun95, EB04, Fox90, FG96, Gur90, GW90, GW91, Ham95, Hol03, Hou92, KH91, KH92, KC92, Kyr95, Lin92, LLAK05, Mah96, Mit88a, Moo03, MP96, MG94, NSW93, Nør00, OC88a, Øst93c, PH09, Par96, PG02, PKB04a, Pre95, QW05, Ram88, Ram91, RM91a, RS89, Sab94, Sah02, Sei09, Sha04, Shu93, SC93a, SS91b, SS91a, Tun89a, Wit96c, Wu90, Zen91, vAK92b, van95, Bir92, Fra97, Kuh89, Sof97, DS94]. Programs [BM90, BA86, CS97, DDPA96, DPDD96, Knu94, Roy06, Sen92, SS92]. Project [RM91b, RM91a, BW91, PG02]. Projekten [Sch92]. proofs [SBR94]. Propaganda [Lee94]. Prototyping [vDHK96, Cze90]. pseudocode [Roy06]. publication [Thi03]. Python [Zuk97].

Quality [SMEN96]. quantum [QW05]. quaternion [AAK01]. query [Cze90]. Queue [Pap90].

R [Sar06, Fil08, GB06, Lei02b, Rac12]. R/Sweave [GB06]. Rambutan [Sah02]. rapid [Cze90]. Readable [BM90]. Real [CH95b]. ReDuX [Mor94]. Refer [Ano10c]. Reference [Chi93]. reflections [Sei09]. Regular [Ham88, VHC88]. Reimagining [PH09]. Reject [Hüs06]. Relational [JAC94]. Remembering [Bri96e]. report [GW90, GW91]. Reports [Lei02a, Pas17]. Representations [RF97]. Reproducible [LR03, SDDD12, MR09]. Requirements [Nør99b, Nør00]. Research [SDDD12, LR03, MR09, OC88a]. Resources [Ska98c]. Response [Syr99]. Restricted [Hüs06]. Retargetable [FH95]. Reusability [Pap90]. Reusable [Han97]. Reuse [CS97, DN07, Hen97, Sit96, CS96]. Reverse [Kna96]. Review [Bir92, Thi89, Dug93]. Revisited [Han94]. roar [NTW91]. Role [BC87, BC90b]. Rome [Gia89]. Routines [Don01a, Don01b]. RStudio [Rac12]. Runoff [Ano10c].

SAC-2 [Kre89]. Sample [Hüs06]. sampling [BK86]. SAS [LH07]. SASweave [LH07]. SCANTEX [Sew87b]. SCHEME [KH91, KH92, Nør99a]. SchemeTEX [Ram88]. School [SMEN96]. Science [Joy94, Ros96]. Sciences [IEE90]. Scientific [Pas17]. Scotland [SMEN96]. Scribus [Ano10a]. Scripting [Ska98c]. Search [Pfa04]. Seattle [Ost93b]. Selected [Knu00, Knu03a, Knu03b, BW91]. Selective [Zuk97]. Self [Gur90]. semantics [Ruc05]. semiliterate [Lum98]. September [FG96]. Set [Hüs06, SM93, SM96]. setting [Knu82, Knu91]. Seventh [BN93]. SGML [Cov98, Kim97a, Kim97b, MGCR96, MG94, Par96, RW96, SM93, SM96]. SGML-based [MG94, MGCR96]. SGML-Lite [MGCR96]. SGML/XML [Cov98]. SIAM [ACM93]. SIGACT [Ano94]. SIGACT-SIGPLAN [Ano94]. SIGCSE [Joy94]. signal [Sla90]. SIGPLAN [Ano94]. Simple [Bri93, Brixx, Ram88, Ram96, SM96, QW05]. Simplified [Ram94b]. Situated [MM94]. skills [Dun95]. Skribe [GS05]. Skribilo [Ano17]. Smalltalk [Øst93c, Ost95, RS89]. Software [AMS97, Ano10a, Ano10b, BDM97, BC90d, DN07, Han97, Kre89, Pur95, SMEN96, Sit96, BC90b, Car95, EB04, Ham95, Sew87a, Sch92]. Software-Projekten [Sch92]. Solutions [BGG^{+}94]. Solver [Don01b, Ram96]. solving [Dun95]. Some [NTW91]. Source [BM86, Lio96, WB89, Bzy95, Pag07, Ska98a]. sources [Bro88b]. Spaniel [Big99, HS98]. Sparse [Bri96b, Bri96c]. Specification [vDHK96]. specifications [Joh96, Sen92]. Specifying [RF97]. SRIG [Pur95]. SRIG-ET [Pur95]. Stability [SS92]. Standard [Pla92]. Standards [IEE96]. Stanford [Knu93b, Knu93c]. Stat [Lum98]. Statistical [Lei02a, LR03, RL03]. statistics [NTW91]. STOL [KH92, KH91]. Structured [Knu83a, KL93, Tun89b, PG02, SBR94, SW96, Tun89a]. Structures [Bet05]. Student [Hur96]. Students [MM94]. study [Pep91]. Style [Chi92b, OC88b, Oma88, OC90b, OC88a]. Suggestions [Ano98]. supercomputer [ST90]. supplement [Ruc15]. support [Ram88]. Suprenum [ST90]. Survival [Lum98]. Sweave [Bet05, Fil08, GB06, Gor08, Kuh06, Lei02a, Lei02b, Lei02c, Lei03a, Lei03b, Lei03c, Lun11, Rac12, Sar06, MR09]. SWEB [SM93, SM96]. Sydney [Ros96]. Symbolic [CM92, Gia89]. Symposium [ACM93, Ano94, Gia89, Joy94, FG96, Mah96]. syntax [Ruc05]. synthesis [Sab94]. System [AO90, HLR97, IEE90, Knu83a, KL93, SP92, SC93b, WB89, Bec86, Car95, GW90, GW91, Pag07, Sew87a, Shu93, SC93a, Wu90, Zen91]. Systems [RB98, Cze90, Gur90].

Tabular [Mit88b]. Tabular- [Mit88b]. Tag [SM93, SM96]. tailored [Gur90]. Taxonomic [Oma88]. Taxonomy [OC88b]. Tbl [Ano10c]. Tcl [USE97]. Tcl/Tk [USE97]. Teach [SC94]. Teaching [CDL95, MM94, NTW91]. Team [RM91b, RM91a]. tears [NSW93]. Technical [Joy94]. Techniques [Han97, LLAK05, MP96, NBC92]. Technology [IEE96, Par96]. Teilrealisierung [Fra97]. testing [Bal89]. TEX [Sof97]. TexiWeb [Rob13]. Text [BM86, Thi89, Lei03b, Sab94]. Theme [KC02]. Theme-Based [KC02]. Third [IEE90, IEE94]. Thirty [Chi10]. thoughts [NTW91]. Time [Hüs06, NSW93]. Tk [USE97]. Tool [BA86, Bri93, Brixx, vAK92a, Bro88b, vAK92b]. Tools [Ano98, BC87, DS97, KC92, Lee94, Ram92, Ram91, ST90]. Tour [Han94]. toxml [Mac98]. Transactions [Jac87, VJW87]. Translating [Mil94]. Trap [Ram94a]. Trap-Based [Ram94a]. Trees [Pfa04]. Troff [Ano10c, Ano10a, Ano10c]. TUG [BP93]. tuneup [Knu21]. Tutorial [Wil00, Ska98b]. Twenty [IEE90, Joy94]. Twenty-fifth [Joy94]. Twenty-Third [IEE90]. Tycho [HLR97]. Typefaces [Knu86c]. types [BG87]. typesetter [Coa98a]. typesetter-independent [Coa98a]. Typesetting [Ano10a]. Typographic [Oma88, OC90b]. Typography [BM90, Knu99]. Typset [Ano10c].

UIMS [Aki92]. UK [SMEN96]. Ultimate [Ano17]. UML [Moo03]. Understanding [Ruc05]. Undo [Zuk97]. University [NBC92, Pur95, Ros96, SMEN96]. UNIX [Lio96]. Unparsing [Ram98]. update [Cze90]. USA [Ost93b, Sit96]. use [Fil08]. User [BN93, DN07, Don01a, Don01b, HLR97, Lei02c, MC91, Wil92, PG02]. Using [Bet05, Cze90, GB06, Gor08, JJ97, Kim97b, Lei02a, LH07, LLAK05, Ost95, Pas17, RW96, SC94, Thi86, Wit96c, Wit96a, Wit96b, Kyr95, Øst93c, Par96, Wit95]. Utility [Thi89].

Validation [RB98]. VAMP [vA90, vAK92a, vAK92b]. Variant [Thi86]. Version [Brixx, KL93]. via [Fil08]. View [Chi92a, Chi92c]. Viewpoint [CS97]. views [JAC94]. Vignettes [Lei03a, Lei03c]. Visualization [BDM97]. volumes [Ruc15]. Vorschläge [Sch92]. vs [Kro90].

WA [Ost93b]. Weave.jl [Pas17]. Weaving [Ram89, Sew89, VR89, Van90a]. WEB [SD95, MC91, Chi93, Chi92b, Kre89, SS92, VR89, Van90a, Nør99a, Thi86, BK86, BKM86, Knu82]. Web-mode [MC91, Chi93]. WEB/Modula [Kre89]. WEB/Modula-2 [Kre89]. Webless [Fox90]. Werkzeuge [SD95]. Werkzeuges [Kuh89]. Wide [Nør99a]. without [NSW93]. Witt [Han95]. Wizard [Spi91]. Words [VHG87, BKM86]. Work [RW96, Sei09]. Workshop [BN93, IEE94, Ost93b, USE97, Lin92]. World [Mau95, Nør99a]. Write [DDPA96, DPDD96]. Writing [Sof97, Vee17, Goo07]. wysiwyg [Wu90, GW90, GW91].

Xindy [Ano10c]. XLISP [Lum98]. XLISP-Stat [Lum98]. XML [Coa98a, Coa98b, Cov98, Dwe00]. XP [EB04].

Yacc [SM96]. Yacc/Lex [SM96]. years [Chi10].

Z [BN93, Nor93, Sen92]. Zealand [Pur95]. zum [Kuh89].

References

Atlamazoglou:2001:ALP

[AAK01] P. E. Atlamazoglou, H. T. Anastassiu, and D. I. Kaklamani. Application of literate-programming principles for the description of a FOR-TRAN 90 extension to quaternion arithmetic. IEEE Antennas and Propagation Magazine, 43(4):104– 114, August 2001. CODEN IAP-MEZ. ISSN 1045-9243.

ACM:1993:PFA

[ACM93] Proceedings of the Fourth Annual ACM-SIAM Symposium on Discrete Algorithms. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89871-313-7. LCCN QA76.9.A43 A34 1993.

Advogato:2000:IDK

[Adv00] Advogato. Interview with Donald E. Knuth. World-Wide Web document, January 25, 2000. URL http:/

                                                /www.advogato.org/article/28.
                                                html.

Appelt:1986:MCW

[AH86] W. Appelt and K. Horn. Multiple changefiles in WEB. TUGboat, 7(1): 20, March 1986. ISSN 0896-3207.

Aitken:1996:AGJ

[Ait96] Gary Aitken. Automatically generating Java documentation: javadoc and the doc comment. Dr. Dobb’s Journal of Software Tools, 21(7): 42, 44, 46–49, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Akiguchi:1992:AUC

[Aki92] C. Akiguchi. Application of UIMS to CASE environment. Journal of the Information Processing Society of Japan = Joho Shori, 33(11):1314– 1323, November 1992. CODEN JOSHA4. ISSN 0447-8053.

Al-Maati:1997:ILP

[AMS97] Shereef Abu Al-Maati and William D.

Shoaff. Integrating literate programming and cleanroom software engineering. In Second Australasian Conference on Computer Science Education, pages 99–106. University of Melbourne, Melbourne, Victoria, Australia, July 2–4, 1997.

Anderson:2002:FLP

[And02] H. Anderson. Formalization and ‘literate’ programming. In IEEE, editor, Eighth Asia-Pacific Software Engineering Conference, 4– 7 December, 2001. APSEC 2001, pages 39–44. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1408-1. ISSN 1530-1362. LCCN ????

Anonymous:1994:PAA

[Ano94] Anonymous, editor. Proceedings of 21st Annual ACM SIGACT-SIGPLAN Symposium on Principles of Programming Languages. ACM Press, New York, NY 10036, USA, ???? 1994. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994.

Anonymous:1998:SFL

[Ano98] Anonymous. Suggestions for frameworks for literate programming tools. World-Wide Web document.,

1998. URL http://www.ems.uq.
edu.au/Public/Lists/xml-litprog-
l/framework.html.

Anonymous:2010:FTS

[Ano10a] Anonymous. Free Typesetting Software: LaTEX, Troff, Scribus, Figlet, Lout, Noweb, Freetype, Graphite, Groff. Books LLC, ????, 2010. ISBN 1-157-24708-3. 44 pp. LCCN ????

Anonymous:2010:SDL

[Ano10b] Anonymous. Software Documentation: Literate Programming, Software Documentation, Docbook, Gnu Free Documentation License. Books LLC, ????, 2010. ISBN 1-156-09458-5. 70 pp. LCCN ????

Anonymous:2010:TNJ

[Ano10c] Anonymous. Troff: Nroff, James Clark, Noweb, Typset and Runoff, Mike Lesk, Groff, Pic Language, Refer, Xindy, Tbl, Eqn, Makeindex, Joe Ossanna, Troff Macro. Books LLC, ????, 2010. ISBN 1-156-87498-X. 46 pp. LCCN ????

Anonymous:2017:SUD

[Ano17] Anonymous. Skribilo: The ultimate document programming framework. Web site, 2017.

Avenarius:1990:FLP

[AO90] Adrian Avenarius and Siegfried Oppermann. FWEB: a literate programming system for Fortran 8X. ACM SIGPLAN Notices, 25(1):52– 58, January 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ben-Ari:1986:FTD

[BA86] Mordechai Ben-Ari. FOREET: a tool for design and documentation of Fortran programs. Software—Practice and Experience, 16(10):915–924, 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ballard:1989:DTD

[Bal89] F. Ballard. ‘executable’ documentation: testing the documentation documenting the testing. In Conference proceedings / SIGDOC 89, November 8–10, 1989, Pittsburgh Hilton and Towers, Pittsburgh, Pennsylvania, pages 143–146. ACM Press, New York, NY 10036, USA, November 1989. ISBN 0-89791-337-X. LCCN QA 76.9 D6 S54 1989.

Beck:1987:ERT

[BC87] K. Beck and W. Cunningham. Expanding the role of tools in a literate programming environment. In CASE’87, Boston, Massachusetts, pages 1–5. ACM Press, New York, NY 10036, USA, 1987. ISBN ???? LCCN ???? URL http://c2.

com/doc/case87.html;   http:/
/www.literateprogramming.com/
lpsmalltalk.pdf.

Brown:1990:HLP

[BC90a] M. Brown and B. Czejdo. A hypertext for literate programming. In Selim G. Akl, F. Fiala, and Waldemar W. Koczkodaj, editors, Advances in computing and information–ICCI ’90 / International Conference on Computing and Information, Niagara Falls, Canada, May 23–26, 1990, Proceedings, pages 250–259. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., May 1990. ISBN 3-540-53504-7. LCCN QA75.5 .I578 1990.

Brown:1990:RLP

[BC90b] Marcus Brown and David Cordes.

The role of literate programming in the software engineering lifecycle. In Proceedings of the 28th annual Southeast Regional Conference, pages 101–107. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-356-6. LCCN ????

Brown:1990:IEL

[BC90c] Marcus E. Brown and Bart Childs. An interactive environment for literate programming. Structured Programming, 11(1):11–26, 1990. CO-DEN STPGEM. ISSN 0935-1183.

Brown:1990:LPA

[BC90d] Marcus E. Brown and David Cordes. Literate programming applied to conventional software design. Structured Programming, 11(2):85–98, 1990. CODEN STPGEM. ISSN 0935-1183.

Brown:1990:LPD

[BC90e] Marcus E. Brown and David Cordes. A literate programming design language. In COMPEURO’90: Proceedings of the 1990 IEEE International Conference on Computer Systems and Software Engineering, May 8–10, 1990, Tel-Aviv, Israel, pages 548–549. IEEE CS Press, Los Alamitos, CA, USA, 1990. ISBN 0-8186-2041-2. LCCN QA76.9.S88 C66 1990.

Baecker:1997:SVD

[BDM97] Ron Baecker, Chris DiGiano, and Aaron Marcus. Software visualization for debugging. Communications of the Association for Computing Machinery, 40(4):44–54, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Shows typeset literate code produced by methods described in [BM90].

Becker:1986:WSE

[Bec86] Helmut Becker. WEB system extensions. TUGboat, 7(2):109, June 1986. ISSN 0896-3207.

Betebenner:2005:UCS

[Bet05] Damian Betebenner. Using control structures with Sweave. R News: the Newsletter of the R Project, 5 (1):40–44, May 2005. CODEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Bentley:1987:PPA

[BG87] Jon Bentley and David Gries. Programming pearls—abstract data types. Communications of the Association for Computing Machinery, 30(4):284–290, April 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bishop:1992:LPL

[BG92] Judy M. Bishop and Kevin M. Gregson. Literate programming and the LIPED environment. Structured Programming, 13(1):23–34, 1992. CODEN STPGEM. ISSN 0935-1183.

Botsford:1994:PCI

[BGG^{+}94] J. Botsford, A. Gawman, M. Gentleman, E. Kidd, K. Lyons, and J. Slonim, editors. Proceedings. CASCON ’94. Integrated Solutions. Nat. Res. Council Canada, Ottawa, Ont., Canada, 1994.

Bigot:1999:LEC

[Big99] Peter A. Bigot. Letters to the Editor: Comment on Spaniel paper. ACM SIGPLAN Notices, 34(2):42– 45, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [HS98, Syr99].

Birss:1992:RLP

[Bir92] Robert C. Birss. Review of Literate Programming (Knuth, D. E; 1992). Computer, 25(9):126–127, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bentley:1986:PPLa

[BK86] Jon Bentley and with Special Guest Oyster, Donald E. Knuth. Programming pearls: Literate programming: a WEB program for sampling. Communications of the Association for Computing Machinery, 29(5):364– 369, May 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprinted in [Knu92, pp. 137–149].

Bentley:1986:PPLb

[BKM86] Jon Bentley, Donald E. Knuth, and M. Douglas McIlroy. Programming pearls: a literate program: a WEB program for common words. Communications of the Association for Computing Machinery, 29(6):471– 483, June 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=

315653. See another solution to the problem of printing the k most common words of a text file in [VHG87]. Reprinted in [Knu92, 151–177].

Baecker:1986:DPE

[BM86] Ronald Baecker and Aaron Marcus.

Design principles for the enhanced presentation of computer program source text. In Proceedings CHI’86 (Human Factors in Computing Systems), pages 51–58. ACM Press, New York, NY 10036, USA, April 1986.

Baecker:1990:HFT

[BM90] Ronald Baecker and Aaron Marcus.

Human Factors and Typography for More Readable Programs. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-10745-7. xx + 348 pp. LCCN QA76.75 .B34 1990. See [BDM97].

Bowen:1993:ZUW

[BN93] J. P. Bowen and J. E. Nicholls, editors. Z User Workshop London 1992. Proceedings of the Seventh Annual Z User Meeting. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-19818-0. LCCN QA76.73.Z2 Z2 1992.

Bzyl:1993:ALP

[BP93] Włodek Bzyl and Tomasz Przechlewski. An application of literate programming: creating a format for the Bulletin of the Polish TUG. TUGboat, 14(3):296–299, October 1993. ISSN 0896-3207.

Briggs:1993:NSL

[Bri93] Preston Briggs. Nuweb, a simple literate programming tool. cs.rice.

                                        edu:/public/preston, Rice Uni-

versity, Houston, TX, USA, 1993.

Briggs:1996:LPA

[Bri96a] Preston Briggs. Literate programming: Automatic parallelization. ACM SIGPLAN Notices, 31(4):11– 14, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:LPSa

[Bri96b] Preston Briggs. Literate programming: Sparse matrix manipulation. ACM SIGPLAN Notices, 31(8):5–7, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:LPSb

[Bri96c] Preston Briggs. Literate programming: Sparse matrix multiplication. ACM SIGPLAN Notices, 31(11):33– 37, November 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:PLPb

[Bri96d] Preston Briggs. Programming: Literate programming: Automatic parallelization. ACM SIGPLAN Notices, 31(4):11–15, April 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:PLPa

[Bri96e] Preston Briggs. Programming: Literate programming: Remembering. ACM SIGPLAN Notices, 31(1):4–5, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:19xx:NVS

[Brixx] Preston Briggs. Nuweb version 0.87b: a simple literate programming tool. World-wide web document., Electrical Engineering and Computer Science Department, Harvard University, Cambridge, MA, USA, 19xx. URL

http://www.eecs.harvard.edu/~
nr/noweb/nuweb.html.

Brown:1988:IEL

[Bro88a] Marcus E. Brown. An Interactive Environment for Literate Programming. Thesis (Ph.D.), Texas A&M University, College Station, TX, USA, August 1988. ix + 102 pp.

Brown:1988:LPT

[Bro88b] Marcus Edward Brown. The literate programming tool: the sources. Technical report TAMU-88-023, Texas A and M University, Computer Science Dept., College Station, TX, USA, August 1988. (various) pp.

Broy:1991:MPS

[BW91] Manfred Broy and Martin Wirsing. Methods of programming. Selected papers on the CIP-Project, volume 544. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CO-DEN LNCSD9. ISBN 3-540-54576-X (Berlin), 0-387-54576-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 268 pp. LCCN QA76.6 .M4543 1991, QA267.A1 L43 no.544. URL http://link.

springer-ny.com/link/service/
series/0558/tocs/t0544.htm;

                                        http://www.springerlink.com/
                                        openurl.asp?genre=issue&issn=
                                        0302-9743&volume=544.

Bzyl:1995:LPS

[Bzy95] Włodek Bzyl. Literate Plain source is available! TUGboat, 16(3):297– 299, September 1995. ISSN 0896-

                                        3207. URL ftp://ftp.ac.tex.
                                        uk/ctan/tex-archive/web/tweb/
                                        tug95/.

Carter:1995:CHS

[Car95] R. Carter. The CASEtool hypertext system for software development. In Maurer [Mau95], pages 133–138.

Cordes:1991:LPP

[CB91] David Cordes and Marcus Brown. The literate-programming paradigm. Computer, 24(6):52–61, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Childs:1995:TCC

[CDL95] Bart Childs, Deborah Dunn, and William Lively. Teaching CS/1 courses in a literate manner. TUGboat, 16(3):300–309, September 1995. ISSN 0896-3207.

Copeland:1995:ILP

[CH95a] Jeffrey Copeland and Jeffrey Haemer. An introduction to literate programming. RS/Magazine, 4 (1):26–30, January 1995. ISSN 1088-0844, 1061-0030.

Copeland:1995:LPR

[CH95b] Jeffrey Copeland and Jeffrey Haemer. Literate programming: a real example, part 1. RS/Magazine, 4(2):32–35, February 1995. ISSN 1088-0844, 1061-0030.

Copeland:1995:LPE

[CH95c] Jeffrey Copeland and Jeffrey Haemer. Literate programming: An example, part 2. RS/Magazine, 4 (3):31–37, March 1995. ISSN 1088-0844, 1061-0030.

Childs:1992:ELP

[Chi92a] Bart Childs. Errata: Literate Programming, A Practitioner’s View, TUGboat 1, no. 3(3), pp. 261– 268. TUGboat, 13(4):457, December 1992. ISSN 0896-3207. See [Chi92c].

Childs:1992:IWS

[Chi92b] Bart Childs. An introduction to the WEB style of literate program-

ming.  ftp.cs.tamu.edu:/pub/
tex-web/web/docs, Texas A&M

University, College Station, TX, USA, 1992.

Childs:1992:LPP

[Chi92c] Bart Childs. Literate programming, a practitioner’s view. TUGboat, 13 (3):261–268, October 1992. ISSN 0896-3207. See errata [Chi92a].

Childs:1993:GER

[Chi93] Bart Childs. GNU Emacs reference card (with web-mode). ftp.

cs.tamu.edu:/pub/tex-web/web/

docs, Texas A&M University, College Station, TX, USA, 1993.

Childs:2010:TYL

[Chi10] Bart Childs. Thirty years of literate programming and more? TUGboat, 31(2):183–188, 2010. ISSN 0896-3207.

Childs:1992:SCA

[CM92] S. B. Childs and T. J. McGuire. Symbolic computing, automatic programming, and literate programming. In Noye et al. [NBC92], pages 29–48. ISBN 0-86396-172-X. LCCN ????

Coates:1998:FDX

[Coa98a] A. B. Coates. funnelweb.dtd: XML DTD for typesetter-independent funnelweb documents. World-Wide Web document., May 01, 1998.

URL http://www.ems.uq.edu.au/
People/Tony/litprog/xml/dtd/
funnelweb.dtd.

Coates:1998:XLP

[Coa98b] Anthony B. Coates. XML and literate programming. World-Wide Web document., 1998. URL http:

//www.ems.uq.edu.au/Seminars/
XML_LitProg/.

Cover:1998:SXL

[Cov98] Robin Cover. SGML/XML and literate programming. World-Wide Web document., 1998. URL http:/

/www.sil.org/sgml/xmlLitProg.

html. This document includes links to other literate-programming-in-SGML documents and software packages.

Childs:1996:LPD

[CS96] Bart Childs and Johannes Sametinger. Literate programming and documentation reuse. In Sitaraman [Sit96], pages 205–214. ISBN 0-8186-7301-X, 0-8186-7303-6. ISSN 1085-9098. LCCN QA76.76.R47 I587 1996. IEEE order plan catalog number 96TH100015.

Childs:1997:ALP

[CS97] Bart Childs and Johannes Sametinger.

Analysis of literate programs from the viewpoint of reuse. Software — Concepts and Tools, 18(1):35– ??, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Czejdo:1990:UQU

[Cze90] B. Czejdo. Using an ER query and update interface for rapid prototyping of hypertext systems. In IEEE [IEE90], pages 227–236 (vol. 2). 4 vol.

Deransart:1996:HBEa

[DDPA96] P. Deransart, R. Da Silva Bigonha, P. Parot, and M. Andrade. A hypertext based environment to write literate logic programs. In Fuchs and Geske [FG96], pages 247–252. ISBN 3-88457-296-2. ISSN 0170-8120. LCCN ????

Denning:1987:ALP

[Den87] Peter J. Denning. Announcing literate programming. Communications of the Association for Computing Machinery, 30(7):593, July 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dickey:1993:LPA

[Dic93] Lee J. Dickey. Literate programming in APL and APLWEB. APL Quote Quad, 23(4):11–??, June 1, 1993. CODEN APLQD9. ISSN 0163-6006.

Dinmore:2007:LMI

[DN07] M. Dinmore and A. F. Norcio. Literacy for the masses: Integrating software and knowledge reuse for end-user developers through literate programming. In IEEE International Conference on Information Reuse and Integration, 13–15 August, 2007. IRI 2007, pages 455– 460. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 1-4244-1500-4. LCCN ????

Donev:2001:CUNa

[Don01a] Aleksandar Donev. Communicating with the user in numerical routines: Part I: OOP approach in F2x. ACM Fortran Forum, 20(2):11–18, August 2001. ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2001:CUNb

[Don01b] Aleksandar Donev. Communicating with the user in numerical routines: Part II: An iterative linear solver in F95 and F2X. ACM Fortran Forum, 20(2):19–32, August 2001. ISSN 1061-7264 (print), 1931-1311 (electronic).

Deransart:1996:HBEb

[DPDD96] P. Deransart, P. Parot, R. Da Silva Bigonha, and M. A. Da Silva Bigonha. A hypertext based environment to write literate logic programs. In Maher [Mah96], pages 540–?? ISBN 0-262-63173-3. LCCN QA76.63 L831 1996.

Detig:1994:DKL

[DS94] Christine Detig and Joachim Schrod. Donald E. Knuth, Literate Programming. TUGboat, 15(1): 25–27, March 1994. ISSN 0896-3207.

Detig:1997:LPC

[DS97] Christine Detig and Joachim Schrod. Literate Programming: Concepts, Methods, and Tools. ITCP-UK Computer Science Series. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 1-85032-305-4. ???? pp. LCCN ???? US$36.95. Includes disk.

Duggan:1993:LPR

[Dug93] Angus Duggan. Literate programming: a review. LFCS report ECS-LFCS-93-263, University of Edinburgh, Laboratory for Foundation of Computer Science, Edinburgh, Scotland, 1993. 25 pp.

Dunn:1995:LPM

[Dun95] Deborah Lynn Byrum Dunn. Literate programming as a mechanism for improving problem solving skills. Thesis (Ph.D.), Texas A&M University, Department of Computer Science, College Station, TX, USA, 1995. xiii + 268 pp.

Dwelly:2000:JXL

[Dwe00] Andrew Dwelly. Java, XML, and literate programming. Dr. Dobb’s Journal of Software Tools, 25(2): 62, 64–66, 68, February 2000. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.com/
ftp/2000/2000_02/litjava.txt;
http://www.ddj.com/ftp/2000/
2000_02/litjava.zip.

Eckstein:2004:EPA

[EB04] Jutta Eckstein and Hubert Baumeister, editors. Extreme programming and agile processes in software engineering: 5th international conference, XP 2004, Garmisch-Partenkirchen, Germany, June 6– 10, 2004: proceedings, volume 3092 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-22137-9 (paperback). LCCN QA76.76.D47 X62 2004. URL

http://www.springerlink.com/
openurl.asp?genre=issue&issn=
0302-9743&volume=3092.

Fuchs:1996:LPJ

[FG96] Norbert E. Fuchs and Ulrich Geske, editors. Logic programming: Joint international conference and symposium — September 2–6, 1996, Bonn, Germany, volume 296 of GMD-Studien. GMD-Forschungszentrum Informationstechnik, Sankt Augustin, Germany, 1996. ISBN 3-88457-296-2. ISSN 0170-8120. LCCN ????

Fraser:1995:RCC

[FH95] Chris W. Fraser and David R. Hanson. A Retargetable C Compiler: Design and Implementation. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1995. ISBN 0-8053-1670-1. xv + 564 pp. LCCN QA76.73.C15F75 1995. US$55.95.

URL https://www.pearson.com/
us/higher-education/program/
Hanson-Retargetable-C-Compiler-
A-Design-and-Implementation/

PGM166351.html. From the authors’ announcement: “… this book is an example of a ‘literate program.’ Like TEX: The Program by D. E. Knuth (Addison-Wesley, 1986), this book is lcc’s source code and the prose that describes it. The code is presented in the order that best suits understanding, not in the order dictated by the C programming language. The source code that appears on the diskette is extracted automatically from the book’s text files. … The distribution is available via ‘anony-

mous’ ftp from ftp.cs.princeton.
edu (128.112.152.13) in the di-

rectory pub/lcc. … Additional information about lcc and about our book is available on the WWW at

URL http://www.cs.princeton.
edu/software/lcc.”.

Filzmoser:2008:LRS

[Fil08] Peter Filzmoser. LaTEX and R via Sweave: An example document how to use Sweave. Technical report, Vienna University of Technology, Wiedner Haupstraße 8-10, A-1040 Wien, Austria, March 12, 2008. 6 pp. URL http:

//www.statistik.tuwien.ac.at/
public/filz/students/SweaveExa.
pdf.

Fox:1990:WLP

[Fox90] Jim Fox. Webless literate programming. TUGboat, 11(4):511– 513, November 1990. ISSN 0896-3207.

Franosch:1997:KTI

[Fra97] Heike Franosch. Konzeption und Teilrealisierung einer interaktiven Programmierumgebung für die Lehrerausbildung basierend auf der Idee des Literate-Programming. Diplomarbeit, Universität Stuttgart, Stuttgart, Germany, 1997. 165 pp.

Garbade:2006:URS

[GB06] Sven Garbade and Peter Burgard. Using R/Sweave in everyday clinical practice. R News: the Newsletter of the R Project, 6(2):26–31, May 2006. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-
                                        project.org/doc/Rnews/.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4–8, 1988: proceedings, volume 358 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Goodliffe:2007:CCP

[Goo07] Pete Goodliffe. Code craft: the practice of writing excellent code. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-119-0. LCCN QA76.6 .G656 2007.

Gorjanc:2008:USL

[Gor08] Gregor Gorjanc. Using Sweave with LyX. R News: the Newsletter of the R Project, 8(1):2–9, May 2008. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-
                                        project.org/doc/Rnews/.

Guntermann:1986:WAC

[GS86] Klaus Guntermann and Joachim Schrod. WEB adapted to C. TUGboat, 7(3):134–137, October 1986. ISSN 0896-3207.

Gallesio:2005:SFA

[GS05] Erick Gallesio and Manuel Serrano. Skribe: a functional authoring language. Journal of Functional Programming, 15(5): 751–770, September 2005. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL  https://www.cambridge.
org/core/product/689EF8960521587B23203A56C46D2592.

Gurari:1990:STS

[Gur90] Eitan Gurari. Self tailored systems for literate programming. Technical research report OSU-CISRC-2/90-TR4, Ohio State University, Computer and Information Science Research Center, Columbus, OH, USA, 1990. 33 pp.

Gurari:1994:TLD

[Gur94] Eitan M. Gurari. TEX and LaTEX:

Drawing and Literate Programming. McGraw-Hill programming tools for scientists and engineers. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025208-4 (book), 0-07-911616-7 (book + diskette). xiv + 310 pp. LCCN Z253.4.T47G87, 686.2’2544536-dc20, 93-8603 CIP. US$34.95.

Gurari:1990:WLP

[GW90] Eitan Gurari and Jesse Wu. A WYSIWYG literate programming system: a preliminary report. Technical research report OSU-CISRC-7/90-TR17, Ohio State University, Computer and Information Science Research Center, Columbus, OH, USA, 1990. 15 pp.

Gurari:1991:WLP

[GW91] E. M. Gurari and J. Wu. A WYSIWYG literate programming system (preliminary report). In 1991 ACM Computer Science Conference: March 5–7, 1991, San Antonio Convention Center, San Antonio, Texas: Proceedings: Preparing for the 21st Century, pages 94–104. ACM Press, New York, NY 10036, USA, March 1991. ISBN 0-89791-382-5. LCCN QA75.5 .A137 1991.

Hamilton:1988:LPE

[Ham88] Eric Hamilton. Literate programming—expanding generalized regular expressions. Communications of the Association for Computing Machinery, 31(12):1376–1385, December 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamer:1995:LPS

[Ham95] J. Hamer. Literate programming: a software engineering perspective. In Purvis [Pur95], pages 282– 288. ISBN 0-8186-5870-3. LCCN QA76.758 .S6447 1994.

Hansen:1994:MLD

[Han94] Per Brinch Hansen. Multiplelength division revisited: a tour of the minefield. Software—Practice and Experience, 24(6):579–601, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). This paper derives an algorithm for division of long integers, and implements it as a literate program, although without identifier cross-references. See comment about another division algorithm [Han95].

Hansen:1995:LDA

[Han95] Per Brinch Hansen. The long division algorithm of Linger, Mills and Witt. Software—Practice and Experience, 25(1):109, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Han94].

Hanson:1997:CII

[Han97] David R. Hanson. C Interfaces and Implementations: Techniques for Creating Reusable Software. Addison-Wesley Professional Computing Series. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-49841-3. xvii + 519 pp. LCCN QA76.73.C15H37 1997. US$37.95.

URL http://www.cs.princeton.
edu/software/cii/.

Hendseth:1997:LPC

[Hen97] Sverre Hendseth. Literate programming and code reuse. Dr. Dobb’s Journal of Software Tools, 22(6):18– 20, 22, 24, 97, 98, June 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Hylands:1997:TUI

[HLR97] Christopher Hylands, Edward A.

Lee, and H. John Reekie. The Tycho user interface system. In USENIX [USE97], pages 149–??

Holub:1990:CDC

[Hol90] Allen I. Holub. Compiler Design in C. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-155045-4, 0-13-155151-5 (paperback). xviii + 924 pp. LCCN QA76.76.C65 H65

1990. URL http://www.holub.

                                        com/software/compiler.design.
                                        in.c.html.

Holmes:2003:CPP

[Hol03] N. Holmes. The case for perspicuous programming. Computer, 36(4):104, 102–103, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4104.htm;
                                         http://csdl.computer.org/dl/
                                        mags/co/2003/04/r4104.pdf.

Houser:1992:LLP

[Hou92] Christopher Driessen Houser. A literate Lisp programming environment. Thesis (M.S.), Department of Computer Science, University of California, Los Angeles, Los Angeles, CA, USA, 1992. viii + 93 pp. Typescript (photocopy).

Hatzigeorgiu:1998:LPS

[HS98] Nick Hatzigeorgiu and Apostolos Syropoulos. Literate programming and the “Spaniel” model. ACM SIGPLAN Notices, 33(12):52–56, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [Big99, Syr99].

Hurst:1996:LPA

[Hur96] A. J. Hurst. Literate programming as an aid to marking student assignments. In Rosenberg [Ros96], pages 280–286. ISBN 0-89791-845-2. LCCN ???? ACM Order no. 457963.

Husing:2006:ARA

[Hüs06] Johannes Hüsing. An accept-and-reject algorithm to sample from a set of permutations restricted by a time constraint. Journal of Statistical Software, 16(6):1–17, May 2006. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.org/
v16/i06.

Hyman:1990:LC

[Hym90] Marco S. Hyman. Literate C++.

Computer Language Magazine, 7(7): 67–68, 70, 72, 74–77, 79, July 1990. CODEN COMLEF. ISSN 0749-2839.

IEEE:1990:PTA

[IEE90] IEEE, editor. Proceedings of the Twenty-Third Annual Hawaii International Conference on System Sciences. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. 4 vol.

IEEE:1994:PIT

[IEE94] IEEE, editor. Proceedings. IEEE Third Workshop on Program Comprehension. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5647-6. LCCN QA76.751 .I38 1994. IEEE Catalog No. 94TH06767.

IEEE:1996:CPE

[IEE96] IEEE, editor. COMPASS ’96:

proceedings of the Eleventh Annual Conference on Computer Assurance, June 17–21, 1996, National Institute of Standards and Technology, Gaithersburg, Maryland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3390-X. LCCN QA 76.76 R44 C668 1996. IEEE Catalog No. 96CH35960.

Jackson:1987:LPP

[Jac87] Michael A. Jackson. Literate programming—processing transactions. Communications of the Association for Computing Machinery, 30(12):1000–1010, December 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:1994:RVP

[JAC94] Tim Jones, Warwick Allison, and David Carrington. Relational views for program comprehension. In IEEE [IEE94], pages 136–144. ISBN 0-8186-5647-6. LCCN QA76.751 .I38 1994. IEEE catalog number 94TH06767.

Johnson:1997:LPU

[JJ97] Andrew L. Johnson and Brad C. Johnson. Literate programming using Noweb. Linux Journal, 42:64– 69, October 1997. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL ftp:/

                                        /ftp.ssc.com/pub/lj/listings/
                                        issue42/2188.tgz.

Johnson:1996:LS

[Joh96] C. W. Johnson. Literate specifications. Software Engineering Journal, 11(4):225–237, July 1996. CO-DEN SEJOED. ISSN 0268-6961.

Joyce:1994:PTS

[Joy94] D. Joyce, editor. The Papers of the Twenty-fifth SIGCSE Technical Symposium on Computer Science Education: Phoenix, Arizona, March 10–11, 1994, volume 26(1) of SIGCSE Bulletin. ACM Press, New York, NY 10036, USA, 1994. CO-DEN SIGSD3. ISBN 0-89791-646-8. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.A1 A1125 v.26:1.

Kortright:1992:CCT

[KC92] E. Kortright and D. Cordes. Cnest and Cscope: Tools for the literate programming environment. In Proceedings / IEEE Southeastcon ’92, April 12–15, 1992, Birmingham, Alabama, pages 604–609 (vol. 2). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 1992. ISBN 0-7803-0494-2. LCCN TK5 .I32 1992 v.1-2. 2 vol.

Kacofegitis:2002:TBL

[KC02] Andreas Kacofegitis and Neville Churcher. Theme-based literate programming. In Paul Strooper, editor, Ninth Asia-Pacific Software Engineering Conference, APSEC 2002: 4–6 December 2002, Gold Coast, Australia: Proceedings, pages 549–557. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1850-8. LCCN

???? URL http://www.gbv.de/
dms/bowker/toc/9780769518503.

IEEE Computer Society order number PR01850.

Kennedy:1988:TAC

[Ken88] David Kennedy. TEX adapted to CWEB. TUGboat, 9(2):124–125, August 1988. ISSN 0896-3207.

Kobler:1991:SLP

[KH91] Daniel Kobler and Daniel Hernandez. StoL — a literate programming in SCHEME. Forschungsberichte Künstliche Intelligenz DDB DBF DKb 92/1447, Technische Universität München, Institut für Informatik, München, Germany, 1991. 20 pp.

Kobler:1992:SLP

[KH92] Daniel Kobler and Daniel Hernández. STOL — literate programming in SCHEME. ACM SIGPLAN Lisp Pointers, 5(4):21–30, October 1992.

Kimber:1997:ALP

[Kim97a] W. Eliot Kimber. An approach to literate programming with SGML architectures. World-Wide Web document., July 18, 1997. URL

                                        http://www.isogen.com/papers/
                                        litprogarch/litprogarch.html.

Kimber:1997:USA

[Kim97b] W. Eliot Kimber. Using SGML architectures and DSSSL to do literate programming. World-Wide Web document., July 20, 1997.

                                        URL http://www.sil.org/sgml/
                                        kimberDSSSLLitProg.html.

Knuth:1993:CSS

[KL93] Donald E. Knuth and Silvio Levy. The CWEB System of Structured Documentation, Version 3.0. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-57569-8. 226 pp. LCCN QA76.9.D3 K6 1993.

Knasmueller:1996:RLP

[Kna96] M. Knasmueller. Reverse literate programming. In Samson et al.

[SMEN96], pages 97–105. ISBN 1-899796-02-9 (??invalid ISBN??). LCCN QA76.76.Q35 S64 1996.

Knuth:1982:FPG

[Knu82] Donald Knuth. Fixed-point glue setting—an example of WEB. TUGboat, 3(1):10, March 1982. ISSN 0896-3207.

Knuth:1983:WSS

[Knu83a] Donald E. Knuth. The WEB system of structured documentation. Stanford Computer Science Report CS980, Stanford University, Stanford, CA, September 1983.

Knuth:1983:LP

[Knu83b] Donald E. Knuth. Literate programming. Technical report STAN-CS-83-981, Stanford University, Department of Computer Science, 1983.

Knuth:1984:LP

[Knu84] Donald E. Knuth. Literate programming. The Computer Journal, 27(2):97–111, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Knuth:1986:MP

[Knu86a] Donald E. Knuth. : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:1986:TP

[Knu86b] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986.

Knuth:1986:CMT

[Knu86c] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:1991:FPG

[Knu91] Donald Knuth. Fixed-point glue setting: Errata. TUGboat, 12(2):313, June 1991. ISSN 0896-3207.

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6.K644. URL

http://csli-www.stanford.edu/
publications/literate.html.

Knuth:1993:LM

[Knu93a] Donald E. Knuth. The literate mathematician. BBC Radio 5’s series Maths Miscellany, produced by John Jaworski and Giselle Corbett (first broadcast February 21, 1993), 30 min. Cassettes available from BBC OUPC, Walton Hall, Milton Keynes, MK7 6BH, England., February 1993.

Knuth:1993:SGPa

[Knu93b] Donald E. Knuth. Stanford Graph-Base: a platform for combinatorial algorithms. In ACM-SIAM-DA4 [ACM93], pages 41–43. ISBN 0-89871-313-7. LCCN QA76.9.A43 A34 1993.

Knuth:1993:SGPb

[Knu93c] Donald E. Knuth. The Stanford GraphBase: a Platform for Combinatorial Computing. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-201-54275-7. vii + 576 pp. LCCN QA164 .K6 1993. US$45.25. [From the publisher]: … represents Knuth’s final preparation for Volume 4 of The Art of Computer Programming. Through the use of about 30 examples, the book demonstrates the art of literate programming. Each example is a programmatic essay, a short story that can be read by human beings, as well as read and interpreted by machines. In these essays/programs, Knuth makes new contributions to the exposition of several important algorithms and data structures.

Knuth:1994:MIL

[Knu94] Donald E. Knuth. Mini-indexes for literate programs. Software — Concepts and Tools, 15(1):2–11, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Knuth:1999:DT

[Knu99] Donald E. Knuth. Digital Typography, volume 78 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 1999. ISBN 1-57586-010-4 (paperback), 1-57586-011-2 (hardcover). xvi + 685 pp. LCCN Z249.3.K59 1998. US$90.00 (cloth), US$39.95 (paperback).

Knuth:2000:SPA

[Knu00] Donald E. Knuth. Selected Papers on Analysis of Algorithms, volume 102 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2000. ISBN 1-57586-212-3 (paperback), 1-57586-211-5 (cloth). xvi + 621 pp. LCCN QA9.58 .K65 2000.

Knuth:2003:SPC

[Knu03a] Donald E. Knuth. Selected Papers on Computer Languages. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-381-2 (hardback), 1-57586-382-0 (paperback). xvi + 594 pp. LCCN QA76.7 .K63 2002. US$75.00 (hardback), US$35.00 (paperback).

Knuth:2003:SPD

[Knu03b] Donald E. Knuth. Selected Papers on Discrete Mathematics, volume 106 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-248-4. xvi + 812 pp. LCCN QA39.3 .K59 2001. US$40.00.

Knuth:2021:TT

[Knu21] Donald Knuth. The TEX tuneup of 2021. TUGboat, 42(1):1–4 (check??), February 2021. ISSN

                                        0896-3207. URL https://tug.org/
                                        TUGboat/tb42-1/tb130knuth-tuneup21.
                                        pdf.

Kredel:1989:SDC

[Kre89] H. Kredel. Software development for computer algebra or from ALDES/ SAC-2 to WEB/Modula-2. In Gianni [Gia89], pages 447–455. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Krommes:1990:FKV

[Kro90] John Krommes. FWEB (Krommes)

vs. FWEB (Avenarius and Oppermann). TEXhax, 90(19), February 1990.

Kuhlmann:1989:EIW

[Kuh89] Karl-Wilhelm Kuhlmann. Entwurf und Implementierung eines Werkzeuges zum Literate Programming. Diplomarbeit, Universität Bielefeld, Fakultät für Wirtschaftswiss, Bielefeld, Germany, 1989. 115 pp.

Kuhn:2006:SOD

[Kuh06] Max Kuhn. Sweave and the Open Document Format — the odfWeave package. R News: the Newsletter of the R Project, 6(4):2–8, October 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Kyriazopoulos:1995:ILP

[Kyr95] Evan Kyriazopoulos. Introduction to literate programming using noweb. SunOpsis: The Developer Support Centre Newsletter, 4(3), July–November 1995. URL

http://www.sun.ca/newsletter/
SunOpsisVol4.No3/SunOpsisVol4.

No3.html#12256. Available elec-

tronically at http://www.sun.ca/
newsletter/SunOpsisVol4.No3/
SunOpsisVol4.No3.html#12256.

Lecarme:1985:LP

[Lec85] O. Lecarme. Literate programming.

ACM Computing Reviews, 26(1):75, 1985.

Lee:1994:LPP

[Lee94] C. Lee. Literate programming — propaganda and tools. Web docu-

ment, 1994. URL http://www.cs.
cmu.edu/~vaschelp/Programming/
Literate/literate.html.

Leisch:2002:SDG

[Lei02a] F. Leisch. Sweave: Dynamic generation of statistical reports using literate data analysis. In Wolfgang Härdle and Bernd Rönz, editors, Compstat 2002: Proceedings in Computational Statistics: 15th symposium held in Berlin, Germany, 2002, pages 565–580. Physika-Verlag Heidelberg, Heidelberg, Germany, 2002. ISBN 3-7908-1517-9. LCCN ???? URL

http://www.ci.tuwien.ac.at/~
leisch/Sweave;   http://www.
quantlet.de/scripts/compstat2002_
wh/paper/full/S_08_leisch.pdf;
 http://www.statistik.lmu.de/
~leisch/Sweave/Sweave-compstat2002.
pdf.

Leisch:2002:SPM

[Lei02b] Friedrich Leisch. Sweave, Part I: Mixing R and LaTEX. R News, 2 (3):28–31, December 2002. URL

http://www.statistik.lmu.de/
~leisch/Sweave/Sweave-Rnews-
2002-3.pdf.

Leisch:2002:SUM

[Lei02c] Friedrich Leisch. Sweave user manual. Technical report, Institut für Statistik und Wahrscheinlichkeitstheorie, Technische Universität Wien, Wiedner Haupstraße 8-10, A-1040 Wien, Austria, October 7, 2002. URL http:/

/biosun1.harvard.edu/courses/
individual/bio271/lectures/L7/
Sweave-manual-20021007.pdf;
http://rss.acs.unt.edu/Rdoc/
library/utils/html/Rtangle.html;
 http://rss.acs.unt.edu/Rdoc/
library/utils/html/RweaveLatex.
html;   http://rss.acs.unt.edu/
Rdoc/library/utils/html/Sweave.
html.

Leisch:2003:PPI

[Lei03a] Friedrich Leisch. Sweave, Part II:

Package vignettes. R News: the Newsletter of the R Project, 3(2): 21–24, October 2003. CODEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Leisch:2003:SBC

[Lei03b] Friedrich Leisch. Sweave and beyond: Computations on text documents. In Kurt Hornik, Friedrich Leisch, and Achim Zeileis, editors, Proceedings of the 3rd International Workshop on Distributed Statistical Computing, Vienna, Austria, 2003, page ?? ????, ????, 2003. ISBN ???? ISSN 1609-395X. LCCN ????

Leisch:2003:SPI

[Lei03c] Friedrich Leisch. Sweave, Part II: Package vignettes. R News, 3(2):21–24, October 2003. URL

http://www.statistik.lmu.de/
~leisch/Sweave/Sweave-Rnews-
2003-2.pdf.

Levy:1987:WAC

[Lev87a] Silvio Levy. WEB adapted to C, another approach. TUGboat, 8(1):12– 13, April 1987. ISSN 0896-3207.

Levy:1987:LPC

[Lev87b] Silvio Levy. Literate programming in C. TEXniques, Publications for the TEX community, 5:125–130, 1987.

Levy:1993:LPC

[Lev93] Silvio Levy. Literate programming and Cweb. Computer Language Magazine, 10(1):67–68, 70, January 1993. CODEN COMLEF. ISSN 0749-2839.

Lenth:2007:SLP

[LH07] Russell V. Lenth and Søren Højsgaard. SASweave: Literate programming using SAS. Journal of Statistical Software, 19(8):1– 20, May 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v19/i08.

Lins:1989:FLA

[Lin89a] Charles Lins. A first look at literate programming. Structured Programming, 10(1):60–62, 1989. CODEN STPGEM. ISSN 0935-1183.

Lins:1989:ILP

[Lin89b] Charles Lins. An introduction to literate programming. Structured Programming, 10(2):107–112, 1989. CODEN STPGEM. ISSN 0935-1183.

Lindenberg:1992:LPW

[Lin92] Norbert Lindenberg. Literate programming workshop, 1992.

                                        URL http://www.desy.de/user/
                                        projects/LitProg/LPW.html.

Lions:1996:LCU

[Lio96] John Lions. Lions’ Commentary on UNIX 6th Edition, with Source Code. Computer classics revisited. Peer-to-Peer Communications, San Jose, CA 95164-0218, USA, 1996. ISBN 1-57398-013-7. 254 pp. LCCN ???? US$29.96; CAN$41.95.

URL http://www.peer-to-peer.
com/catalog/opsrc/lions.html.

With forewords by Dennis M. Ritchie and Ken Thompson. Prefatory notes by Peter H. Salus and Michael Tilson; a Historical Note by Peter H. Salus; and Appreciations by Greg Rose, Mike O’Dell, Berny Goodheart, Peter Collinson, and Peter Reintjes. Originally circulated as two restricted-release volumes: “UNIX Operating System Source Code Level Six”, and “A Commentary on the UNIX Operating System”.

Lymperopoulos:2005:UOO

[LLAK05] D. G. Lymperopoulos, D. Logothetis, P. Atlamazoglou, and D. L. Kaklamani. Using objectoriented and literate-programming techniques for the development of a computational electromagnetics library. IEEE Antennas and Propagation Magazine, 47(3):31–38, June 2005. CODEN IAPMEZ. ISSN 1045-9243.

Leisch:2003:RSR

[LR03] Friedrich Leisch and Anthony J.

Rossini. Reproducible statistical research. Chance: a magazine for people interested in the analysis of data, 16(2):46–50, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lumley:1998:SAX

[Lum98] Thomas Lumley. Survival analysis in XLISP-Stat. A semiliterate program. Journal of Statistical Software, 3(2):1–90, 1998. CO-DEN JSSOBK. ISSN ???? URL

                                        http://www.jstatsoft.org/v03/
                                        i02;   http://www.jstatsoft.
                                        org/v03/i02/code/;   http://
                                        www.jstatsoft.org/v03/i02/coxregdoc.
                                        pdf;   http://www.jstatsoft.
                                        org/v03/i02/coxregdoc.ps.gz;
                                         http://www.jstatsoft.org/v03/
                                        i02/survival.tar.gz;   http:
                                        //www.jstatsoft.org/v03/i02/
                                        updates.

Lundholm:2011:ICC

[Lun11] Michael Lundholm. Implementing the compendium concept with Sweave and DOCSTRIP. The R Journal, 3(2):16–21, December 2011. CODEN ???? ISSN

                                        2073-4859. URL http://journal.
                                        r-project.org/archive/2011-2/
                                        RJournal_2011-2_Lundholm.pdf.

MacKichan:1998:TBE

[Mac98] Barry MacKichan. The toxml back end for noweb. World-Wide Web document., 1998. URL http://

                                        www.halcyon.com/barrym/toxml.
                                        htm.

Maher:1996:LPP

[Mah96] Michael Maher, editor. Logic programming: proceedings of the 1996 joint international conference and symposium on logic programming. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-63173-3. LCCN QA76.63 L831 1996.

Maurer:1995:EMH

[Mau95] H. Maurer, editor. Educational Multimedia and Hypermedia, 1995. Proceedings of ED-MEDIA 95 — World Conference on Educational Multimedia and Hypermedia. Assoc. Advancement of Comput. Educ, Charlottesville, VA, USA, 1995.

Motl:1991:UMG

[MC91] Mark B. Motl and Bart Childs. A user’s manual for GNU Emacs’ web-

mode.  ftp.cs.tamu.edu:/pub/
tex-web/web/docs, Texas A&M

University, College Station, TX, 1991.

Mittelbach:1992:DP

[MDB92] Frank Mittelbach, Denys Duchier, and Johannes Braams. The docstrip program. Technical report, Universität Mainz, Mainz, Germany, 1992. The doc system is a literate programming facility used for much of the LaTEX Project software development.

MoralesGerman:1994:SPE

[MG94] D. Morales-Germán. An SGMLbased programming environment for literate programming. In Botsford et al. [BGG^{+}94], pages 42–49.

Morales-German:1996:SLS

[MGCR96] Daniel Morales-Germán, Donald D. Cowan, and A. Ryman. SGML-Lite — an SGML-based programming environment for literate programming. Technical Report CS-96-6, Department of Computer Science, University of Waterloo, Waterloo, ON, Canada, 1996. 10 pp. URL

http://csgwww.uwaterloo.ca/~
dmg/research/isacc96.ps.gz.

Milligan:1994:TME

[Mil94] W. L. Milligan. Translating M into English. M Computing, 2(4):22– 24, 26, September 1994. CODEN MCPUEF. ISSN 1060-7684.

Mitchell:1988:LP

[Mit88a] Richard Mitchell. Literate programming. Thesis (Ph.D.), Hatfield Polytechnic, Hatfield, West Yorkshire, UK, February 1988. 279 pp.

Mittelbach:1988:NIA

[Mit88b] Frank Mittelbach. A new implementation of the array- and tabularenvironments. TUGboat, 9(3):298– 314, 1988. ISSN 0896-3207.

Meter:1994:EST

[MM94] Glenn Meter and Philip Miller. Engaging students and teaching modern concepts: Literate, situated, object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):329–333, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mehlhorn:1999:LPC

[MN99] Kurt Mehlhorn and Stefan Näher. LEDA: a platform for combinatorial and geometric computing. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-56329-1 (hardcover). xvi + 1018 pp. LCCN QA76.73.C153 M44

                                        1999. URL http://www.loc.gov/
                                        catdir/description/cam0210/99024952.
                                        html;   http://www.loc.gov/catdir/
                                        toc/cam027/99024952.html.

Moon:2003:LPM

[Moo03] Alun Moon. Literate programming meets UML. TUGboat, 24(1):116– 119, 2003. ISSN 0896-3207.

Morin:1994:TRP

[Mor94] Richard Morin. TEX, ReDuX (part II). SunExpert, 5(12):28–31, December 1994. ISSN 1053-9239. Lists sources of TEX, and gives short reviews of several books about TEX, AMS-TEX, LaTEX, and literate programming.

Motl:1990:LPE

[Mot90] Mark Bentley Motl. A Literate Programming Environment Based on an Extensible Editor. Thesis (Ph.D.), Texas A&M University, College Station, TX, USA, December 1990. viii + 110 pp.

Moore:1996:IAL

[MP96] A. P. Moore and C. N. Payne, Jr. Increasing assurance with literate programming techniques. In IEEE [IEE96], pages 187–198. ISBN 0-7803-3390-X. LCCN QA 76.76 R44 C668 1996. IEEE Catalog No. 96CH35960.

Meredith:2009:TRE

[MR09] Evan Meredith and Jeffrey S.

Racine. Towards reproducible econometric research: the Sweave framework. Journal of Applied Econometrics, 24(2):366–374, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Noye:1992:CTA

[NBC92] John Noye, Basil Benjamin, and Lee Colgan, editors. Computational techniques and applications: proceedings of 5th International Computational Techniques and Applications Conference, held at The University of Adelaide, 14– 17 July, 1991, Computational Techniques and Applications 1991. Australian Mathematics Society, Adelaide, South Australia, Australia, 1992. ISBN 0-86396-172-X. LCCN ????

Normington:1993:CZ

[Nor93] G. Normington. Cleanroom and Z. In Bowen and Nicholls [BN93], pages 281–293. ISBN 3-540-19818-0. LCCN QA76.73.Z2 Z2 1992.

Normark:1999:PWW

[Nør99a] Kurt Nørmark. Programming World Wide Web pages in Scheme. ACM SIGPLAN Notices, 34(12):37–47, December 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.cs.auc.
dk/~normark/laml/. Describes a

literate programming environment, LAML (Lisp Abstracted Markup Language), for HTML dynamically generated from Scheme programs.

Normark:1999:REP

[Nør99b] Kurt Nørmark. Requirements for an elucidative programming environment. About to be submitted for publication., 1999.

URL  http://www.cs.auc.dk/~
normark/elucidative-programming/
index.html.

Normark:2000:REP

[Nør00] K. Nørmark. Requirements for an elucidative programming environment. In IEEE, editor, 8th International Workshop on Program Comprehension, 10–11 June, 2000, Limerick, Ireland. Proceedings. IWPC 2000, pages 119–128. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0656-9. LCCN ???? IEEE Computer Society order number PR00656.

Naeve:1993:APT

[NSW93] P. Naeve, B. Strohmeier, and P. Wolf. APL programming without tears to it is time for a change. APL Quote Quad, 24(1):185–189, August 1993. CODEN APLQD9. ISSN 0163-6006. International Conference on APL.

Naeve:1991:HMT

[NTW91] P. Naeve, D. Trenkler, and H. P.

Wolf. How to make the teaching of statistics roar. some thoughts on computer based experiments. Computational Statistics Quarterly, 6(4): 325–353, ???? 1991. CODEN CSQUEM. ISSN 0723-712X.

Oman:1988:PPS

[OC88a] P. W. Oman and C. R. Cook. A paradigm for programming style research. ACM SIGPLAN Notices, 23 (12):69–78, December 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oman:1988:PST

[OC88b] Paul W. Oman and Curtis Cook. A programming style taxonomy. Technical Report 88-60-20, Oregon State University, Corvallis, OR, 1988.

Oman:1990:BPI

[OC90a] Paul W. Oman and Curtis Cook. The book paradigm for improved maintenance. IEEE Software, 7 (1):39–45, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Oman:1990:TSM

[OC90b] Paul W. Oman and Curtis Cook. Typographic style is more than cosmetic. Communications of the Association for Computing Machinery, 33(5):506–520, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oman:1988:TAT

[Oma88] Paul W. Oman. A Taxonomic Analysis of Typographic Programming Style. PhD thesis, Oregon State University, Corvallis, OR, USA, December 1988.

Osterbye:1993:HPD

[Ost93a] K. Osterbye. Hyperstructure program development environments: an experiment with literate programming. In Hypertext in engineering: Workshop: Hypertext ’93 conference — November 1993, Seattle, WA, USA [Ost93b], pages 7–?? ISSN 0908-1216.

Osterbye:1993:HEW

[Ost93b] K. Osterbye, editor. Hypertext in engineering: Workshop: Hypertext ’93 conference — November 1993, Seattle, WA, USA, number 2030 in University of Aalborg Department of Mathematics and Computer Science — Reports — R 1993. Aalborg University, Aalborg, Denmark, 1993. ISSN 0908-1216.

Osterbye:1993:LSP

[Øst93c] Kasper Østerbye. Literate Smalltalk programming using hypertext. Report R 93-2025, University of Aalborg, Institute for Electronic Systems, Dept. of Mathematics and Computer Science (Aalborg universitetscenter. Afdeling for matematik og datalogi), Aalborg, Denmark, August 1993. 29 pp.

Osterbye:1995:LSP

[Ost95] K. Osterbye. Literate Smalltalk programming using hypertext. IEEE Transactions on Software Engineering, 21(2):138–145, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Page:2007:AOS

[Pag07] William S. Page. Axiom: open source computer algebra system. ACM Communications in Computer Algebra, 41(3):114, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Pappas:1990:LPR

[Pap90] T. L. Pappas. Literate programming for reusability: a queue package example. In Proceedings of the Eighth Annual Conference on Ada Technology, pages 500–514. ANCOST, Inc., Atlanta, Georgia, USA, March 1990.

Parker:1996:LPU

[Par96] Chris Parker. Literate programming using SGML and modern hypertext technology. Thesis (M.S.), Department of Computer Science, Worcester Polytechnic Institute, Worcester, MA, USA, 1996. vii, 111 pp.

Pastell:2017:PWJ

[Pas17] Matti Pastell. Weave.jl: Scientific reports using Julia. Journal of Open Source Software, 2 (11):204:1, March 2017. CODEN ???? ISSN 2475-9066. URL

http://joss.theoj.org/papers/
10.21105/joss.00204.

Pepper:1991:LPD

[Pep91] P. Pepper. Literate program derivation: a case study. In Methods of programming. Selected papers on the CIP-Project [BW91], pages 101– 124. CODEN LNCSD9. ISBN 3-540-54576-X (Berlin), 0-387-54576-X (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M4543 1991, QA267.A1 L43 no.544. URL http://link.

springer-ny.com/link/service/
series/0558/tocs/t0544.htm;
http://www.springerlink.com/
openurl.asp?genre=issue&issn=
0302-9743&volume=544.

Pfaff:2004:IBS

[Pfa04] Ben Pfaff. An Introduction to Binary Search Trees and Balanced Trees. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel:

(617) 876-3296, 2004. x + 428 pp.

URL ftp://ftp.gnu.org/avl/.

Patry:2002:EUP

[PG02] G. Patry and P. Girard. Enduser programming in a structured dialogue environment: the GIPSE Project. In IEEE, editor, IEEE Symposia on Human-Centric Computing Languages and Environments: proceedings, September 5-7, 2001, Stresa, Italy, pages 212– 219. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7803-7198-4. LCCN QA76.9.H85.I32 2001. URL http:/

/ieeexplore.ieee.org/servlet/
opac?punumber=7811. IEEE cata-

log number 01TH8587.

Palmer:2009:RLP

[PH09] James Dean Palmer and Eddie Hillenbrand. Reimagining literate programming. In Proceeding of the 24th ACM SIGPLAN conference companion on Object oriented programming systems languages and applications, Orlando, Florida, USA, pages 1007–1014. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-768-0. LCCN ????

Pieterse:2004:CCL

[PKB04a] Vreda Pieterse, Derrick G. Kourie, and Andrew Boake. A case for contemporary literate programming. In Proceedings of the 2004 annual research conference of the South African Institute of Computer Scientists and Information Technologists on IT research in developing countries: Stellenbosch, Western Cape, South Africa, pages 2–9. South African Institute for Computer Scientists and Information Technologists, Cape Town, Republic of South Africa, 2004. ISBN

????  LCCN ????  URL

                                        http://citeseerx.ist.psu.edu/
                                        viewdoc/download?doi=10.1.1.
                                        108.5578&rep=rep1&type=pdf.

Pieterse:2004:LPE

[PKB04b] Vreda Pieterse, Derrick G. Kourie, and Andrew Boake. Literate programming to enhance agile methods. In Eckstein and Baumeister [EB04], pages 250–253. ISBN 3-540-22137-9 (paperback). LCCN QA76.76.D47

                                        X62 2004.  URL http://www.
                                        cs.up.ac.za/cs/vpieterse/pub/
                                        XP2004_LP.pdf.

Jones:1992:IFL

[PL92] Simon L. Peyton Jones and David R. Lester. Implementing functional languages. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-721952-0. xvii + 281 pp. LCCN QA76.62 .P49 1992. US$42.00. Presents literate programs in TEX and Miranda.

Plauger:1992:SCL

[Pla92] P. J. Plauger. The Standard C Library. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-838012-0. xiv + 498 pp. LCCN QA76.73.C15 P563 1991.

Prechelt:1995:CCM

[Pre95] Lutz Prechelt. The CuPit compiler for the MasPar MP-1 and MP-2: a literate programming document. Interner Bericht 95,1, Universität Karlsruhe, Fakultät für Informatik, Karlsruhe, Germany, 1995. 353 pp.

LPT:TB19-2-137

[pt98] LaTEX project team. Default docstrip headers. TUGboat, 19(2):137– 138, June 1998. ISSN 0896-3207.

Purvis:1995:PSE

[Pur95] Martin Purvis, editor. Software Education Conference (SRIG-ET ’94): proceedings, November 22– 25, 1994, University of Otago, New Zealand. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-5870-3. LCCN QA76.758 .S6447 1994.

Quiney:2005:LPQ

[QW05] H. M. Quiney and S. Wilson. Literate programming in quantum chemistry: a simple example. International Journal of Quantum Chemistry, 104(4):430–445, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Racine:2012:RPI

[Rac12] Jeffrey S. Racine. RStudio: A platform-independent IDE for R and Sweave. Journal of Applied Econometrics, 27(1):167–172, January– February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ramsdell:1988:STS

[Ram88] J. D. Ramsdell. SchemeTEX— simple support for literate programming in Lisp. TEXhax, 88(39), April 1988.

Ramsey:1989:WLI

[Ram89] Norman Ramsey. Weaving a language-independent WEB. Communications of the Association for Computing Machinery, 32(9):1051– 1055, September 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramsey:1991:LPTa

[Ram91] Norman Ramsey. Literate programming tools need not be complex. Technical report CS-TR-351-91, Princeton University, Dept. of Computer Science, Princeton, NJ, USA, October 1991. 11 pp.

Ramsey:1992:LPT

[Ram92] Norman Ramsey. Literate-programming tools need not be complex. Re-

port at ftp.cs.princeton.edu in /
reports/1991/351.ps.Z. Software
at ftp.cs.princeton.edu in /pub/
noweb.shar.Z and at bellcore.
com in /pub/norman/noweb.shar.

Z. CS-TR-351-91, Department of Computer Science, Princeton University, August 1992. Submitted to IEEE Software.

Ramsey:1994:CTB

[Ram94a] Norman Ramsey. Correctness of trap-based breakpoint implementations. In Anonymous [Ano94], pages 15–24. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994.

Ramsey:1994:LPS

[Ram94b] Norman Ramsey. Literate programming simplified. IEEE Software, 11(5):97–105, September 1994. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL https://www.cs.tufts.edu/
~nr/pubs/lpsimp.pdf.

Ramsey:1996:SSL

[Ram96] Norman Ramsey. Simple solver for linear equations containing nonlinear operators. Software—Practice and Experience, 26(4):467–487, April 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramsey:1998:UEP

[Ram98] Norman Ramsey. Unparsing expressions with prefix and postfix operators. Software—Practice and Experience, 28(12):1327–1356, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.com/
cgi-bin/abstract?ID=1761;   http:
//www3.interscience.wiley.com/
cgi-bin/fulltext?ID=1761&PLACEBO=
IE.pdf.

Ruys:1998:ELP

[RB98] Theo C. Ruys and Ed Brinksma. Experience with literate programming in the modelling and validation of systems. Lecture Notes in Computer Science, 1384:393–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1384/13840393.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1384/
13840393.pdf.

Ream:2002:LLP

[Rea02a] E. K. Ream. Leo and literate programming. Web document and software., 2002. URL

                                        http://personalpages.tds.net/
                                        %7eedream/design.html.

Ream:2002:LSP

[Rea02b] E. K. Ream. Leo literate editor with outlines. Web document and software., 2002. URL

                                        http://personalpages.tds.net/
                                        %7eedream/intro.html.

Ramsey:1997:SRM

[RF97] Norman Ramsey and Mary F. Fernández. Specifying representations of machine instructions. ACM Transactions on Programming Languages and Systems, 19(3):492–524, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). This article was written using the noweb literate programming system.

Rossini:2003:LSP

[RL03] Anthony J. Rossini and Friedrich Leisch. Literate statistical practice. UW Biostatistics Working Paper Series 194, University of Washington, Seattle, WA, USA,

                                        2003. URL http://www.bepress.
                                        com/uwbiostat/paper194.

Ramsey:1991:LPTb

[RM91a] Norman Ramsey and Carla Marceau. Literate programming on a team project. Technical report CS-TR-302-91, Princeton University, Dept. of Computer Science, Princeton, NJ, USA, February 1991. 11 pp. Published in [RM91b].

Ramsey:1991:LPTc

[RM91b] Norman Ramsey and Carla Marceau. Literate programming on a team project. Software—Practice and Experience, 21(7):677–683, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robbins:2013:TJ

[Rob13] Arnold Robbins. TexiWeb Jr.

Github project., November 2013.

URL https://github.com/arnoldrobbins/
texiwebjr.

Rosenberg:1996:PFA

[Ros96] John Rosenberg, editor. The Proceedings of the First Australasian Conference on Computer Science Education, University of Sydney, Australia, July 3–5, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-845-2. LCCN ???? ACM Order no. 457963.

Rougier:2005:LPC

[Rou05] Jonathan Rougier. Literate programming for creating and maintaining packages. R News: the Newsletter of the R Project, 5(1):35– 39, May 2005. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Roy:2006:DEP

[Roy06] Geoffrey G. Roy. Designing and explaining programs with a literate pseudocode. ACM Journal on Educational Resources in Computing (JERIC), 6(1):1:1–1:??, March 2006. CODEN ???? ISSN 1531-4278.

Reenskaug:1989:ELS

[RS89] Trygve Reenskaug and Anne Lise Skaar. An environment for literate Smalltalk programming. ACM SIGPLAN Notices, 24(10):337–345, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruckert:2005:UMS

[Ruc05] Martin Ruckert. Understanding MP3: syntax, semantics, mathematics, and algorithms. vieweg-it. Vieweg & Son, Braunschweig, Germany, 2005. ISBN 3-528-05905-2 (paperback). xiii + 247 pp. LCCN ML74.4.M6 R83 2005. URL

http://www.gbv.de/dms/bowker/
toc/9783528059057.

Ruckert:2015:MSS

[Ruc15] Martin Ruckert. The MMIX supplement: supplement to The Art of Computer Programming, volumes 1, 2, 3 by Donald E. Knuth. Addison-Wesley, Reading, MA, USA, 2015. ISBN 0-13-399231-4 (paperback), 0-13-399289-6. xxi + 193 pp. LCCN QA76.6 .K64 2005 Suppl. 1. URL

http://mmix.cs.hm.edu/.

Rath:1996:MSW

[RW96] Hans Holger Rath and Hans-Peter Wiedling. Making SGML work: Introducing SGML into an enterprise and using its possibilities in advanced applications. Computer Standards and Interfaces, 18 (1):37–53, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489.

URL http://www.sil.org/sgml/
bib-or.html#rathMacleod.

Sabourin:1994:CCP

[Sab94] Conrad F. Sabourin. Computational character processing: character coding, input, output, synthesis, ordering, conversion, text compression, encryption, display hashing, literate programming: bibliography. Infolingua, Montréal, PQ, Canada, 1994. ISBN 2-921173-18-2. vii + 579 pp. LCCN ????

Saha:2002:RLP

[Sah02] Prasenjit Saha. Rambutan: Literate programming in Java. TUGboat, 23(3/4):304–309, 2002. ISSN 0896-3207.

Sartori:2006:SRL

[Sar06] Nicola Sartori. Sweave = R · LaTEX^{2}. Technical report, Università “Ca’ Foscari” Venezia, Venice, Italy, November 23, 2006. 17

pp. URL http://stat.epfl.ch/
webdav/site/stat/shared/Regression/
EPFL-Sweave-powerdot.pdf.

Simons:1994:LSP

[SBR94] M. Simons, M. Biersack, and R. Raschke. Literate and structured presentation of formal proofs. IFIP Transactions. A. Computer Science and Technology, A-56:61–81, ???? 1994. CODEN ITATEC. ISSN 0926-5473.

Shum:1993:AAP

[SC93a] S. Shum and C. Cook. AOPS:

an abstraction-oriented programming system for literate programming. Software Engineering Journal, 8(3):113–120, May 1993. CO-DEN SEJOED. ISSN 0268-6961.

Shum:1993:AAOb

[SC93b] Stephen Shum and Curtis Cook.

AOPS: an abstraction-oriented programming system for literate programming. Software Engineering Journal, 8(3):113–120, May 1993. CODEN SEJOED. ISSN 0268-6961.

Shum:1994:ULP

[SC94] Stephen Shum and Curtis Cook. Using literate programming to teach good programming practices. In Joyce [Joy94], pages 66–70. CO-DEN SIGSD3. ISBN 0-89791-646-8. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.A1 A1125 v.26:1.

Schulte:1992:KVF

[Sch92] Michael Schulte. Konzeptionelle Vorschläge für den Einsatz von Literate Programming in Software-Projekten. Diplomarbeit, Universität Bielefeld, Bielefeld, Germany, 1992. 88 pp.

Schrod:1995:LPW

[SD95] Joachim Schrod and Christine Detig. Literate Programming — WEB und Co. Methoden und Werkzeuge, Möglichkeiten und Grenzen. International Thomson Publishing, 5101 Madison Road, Cincinnati, OH 45227, USA, 1995. ISBN 3-8266-0176-9. 300 pp. LCCN ???? DM79.00. URL http://www.

ora.de/itp/science/0176/0176.
html.

Schulte:2012:MLC

[SDDD12] Eric Schulte, Dan Davison, Thomas Dye, and Carsten Dominik. A multi-language computing environment for literate programming and reproducible research. Journal of Statistical Software, 46(3):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v46/i03.

Seibel:2009:CWR

[Sei09] Peter Seibel. Coders at work: reflections on the craft of programming. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1948-3. xiii + 617 pp. LCCN HD8039.D37 S45 2009.

Sennett:1992:DCA

[Sen92] C. T. Sennett. Demonstrating the compliance of Ada programs with Z specifications. In Cliff B. Jones, Roger C. Shaw, and Tim Denvir, editors, 5th Refinement Workshop: Proceedings of the 5th Refinement Workshop, organized by BCS-FACS, London, 8–10 January 1992, pages 70–87. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., January 1992. ISBN 3-540-19752-4. LCCN QA76.76.D47 R44 199.

Sewell:1987:HTM

[Sew87a] E. Wayne Sewell. How to MANGLE your software: the WEB system for Modula-2. TUGboat, 8(2):118–122, July 1987. ISSN 0896-3207.

Sewell:1987:TSP

[Sew87b] E. Wayne Sewell. The SCANTEX processor. TUGboat, 8(2):123–128, July 1987. ISSN 0896-3207.

Sewell:1989:WPL

[Sew89] E. Wayne Sewell. Weaving a Program: Literate Programming in WEB. Van Nostrand Reinhold, New York, NY, USA, 1989. ISBN 0-442-31946-0. xx + 556 pp. LCCN QA76.73.W24 S491 1989.

Shankar:2004:ILA

[Sha04] T. R. Shankar. Intersections of literacy acquisition and programming. In J. (John) Hosking, P. (Philip) Cox, et al., editors, 2003 IEEE Symposium on Human Centric Computing Languages and Environments: proceedings: October 28-31, 2003, Auckland, New Zealand, pages 283–284. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7803-8225-

0. LCCN QA76.9.H85.I32 2003; QA76.9.H85.I32. URL http://

ieeexplore.ieee.org/xpl/RecentCon.
jsp?puNumber=8908.

Shum:1993:AAOa

[Shu93] Stephen M. Shum. AOPS: an abstraction oriented programming system for literate programming. Thesis (Ph.D.), Oregon State University, Corvallis, OR, USA, 1993. 153 pp.

Sitaraman:1996:FIC

[Sit96] Murali Sitaraman, editor. Fourth International Conference on Software Reuse: proceedings, April 23– 26, 1996, Orlando, Florida, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7301-X, 0-8186-7303-

6. ISSN 1085-9098. LCCN QA76.76.R47 I587 1996. IEEE order plan catalog number 96TH100015.

Skaller:1998:IDS

[Ska98a] John Skaller. Interscript design and source documentation. World-Wide Web document., 1998. URL http:

//www.triode.net.au/~skaller/
interscript/iscrs.html.

Skaller:1998:IT

[Ska98b] John Skaller. Interscript tutorial. World-Wide Web document.,

1998. URL http://www.triode.
net.au/~skaller/interscript/
iscru.html.

Skaller:1998:MSR

[Ska98c] John Skaller. Maxtal scripting resources: Download. World-Wide Web document., 1998. URL http:

//www.triode.net.au/~skaller/
interscript/iscrdown.html.

Slaney:1990:ISP

[Sla90] Malcolm Slaney. Interactive signal processing documents. IEEE ASSP magazine: a publication of the IEEE Acoustics, Speech, and Signal Processing Society, 7(2):8–20, April 1990. CODEN IAMAEI. ISSN 0740-7467.

Sperberg-McQueen:1993:SST

[SM93] C. M. Sperberg-McQueen. SWEB:

an SGML tag set for literate programming. World-Wide Web document., September 25,

1993. URL http://www.uic.edu/~
cmsmcq/tech/sweb/sweb.html. Re-

vised August 1994; lightly revised March 1995; revised and extended January–March 1996; revised 19 March 1996.

Sperberg-McQueen:1996:SYL

[SM96] C. M. Sperberg-McQueen. A simple yacc/lex processor for Sweb, an SGML tag set for literate programming. World-Wide Web document., February 5,

1996. URL http://www.uic.edu/~
cmsmcq/tech/sweb/swebyacc.html.

Samson:1996:PSQ

[SMEN96] W. B. Samson, I. M. Marshall, and D. G. Edgar-Nevill, editors. Proceedings of the 5th Software Quality Conference: 9 and 10 July 1996, Dudhope Castle, University of Abertay Dundee, Business School, Dundee, Scotland, UK. University of Abertay Dundee, Dundee, Scotland, 1996. ISBN 1-899796-02-9 (??invalid ISBN??). LCCN QA76.76.Q35 S64 1996.

Sofka:1997:WTT

[Sof97] Michael D. Sofka. Writing with TEX, and TEX & LaTEX: Drawing & Literate Programming, by Eitan M. Gurari. TUGboat, 18(1):37–38, March 1997. ISSN 0896-3207.

Sametinger:1992:HSL

[SP92] Johannes Sametinger and Gustav Pomberger. A hypertext system for literate C++ programming. Journal of Object Oriented Programming, 4 (8):24–29, 1992. CODEN JOOPEC. ISSN 0896-8438.

Spivak:1991:LWM

[Spi91] Michael D. Spivak. The L^{A}MS-TEX Wizard’s Manual. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1991.

Smith:1991:ABLb

[SS91a] Lisa M. C. Smith and Mansur H. Samadzadeh. An annotated bibliography of literate programming.

ACM SIGPLAN Notices, 26(1):14– 20, January 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1991:ABLa

[SS91b] Lisa Min-Yi Chen Smith and Mansur H. Samadzadeh. An annotated bibliography of literate programming. Technical report OSU-CS-TR 91-14, Oklahoma State University, Dept. of Computer Science, Stillwater, OK, USA, 1991. 8 pp.

Smith:1992:MCS

[SS92] Lisa M. Smith and Mansur H.

Samadzadeh. Measuring complexity and stability of WEB programs. Structured Programming, 13(1):35– 50, 1992. CODEN STPGEM. ISSN 0935-1183.

Snelting:1990:PTS

[ST90] G. Snelting and C. Thies. Programming tools for the Suprenum supercomputer. In Toulouse ’90. Third International Workshop. Software Engineering and its Applications Proceedings, pages 951–963 (vol. 2). EC2, Nanterre, France, December 1990. 2 vol.

Simons:1996:ALS

[SW96] M. Simons and M. Weber. An approach to literate and structured formal developments. Formal Aspects of Computing, 8(1):86–107, ???? 1996. CODEN FACMES. ISSN 0934-5043.

Swaine:1998:PPL

[Swa98] Michael Swaine. Programming paradigms: Leo and the lizards. Dr. Dobb’s Journal of Software Tools, 23(7):101–103, July 1998. CODEN DDJOEB. ISSN 1044-789X.

Syropoulos:1999:LER

[Syr99] Apostolos Syropoulos. Letters to the Editor: Response to Peter A. Bigot. ACM SIGPLAN Notices, 34(2):46, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [HS98, Big99].

Thimbleby:1984:LPC

[Thi84] Harold W. Thimbleby. Literate programming in C. Technical report, University of York, Department of Computer Science, University of York, Heslington, York YO1 5DD, England, 1984.

Thimbleby:1986:ELP

[Thi86] H. Thimbleby. Experiences of ‘literate programming’ using Cweb (A variant of Knuth’s Web). The Computer Journal, 29(3):201–211, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/201.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/
202.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/203.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/
204.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/205.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/
206.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/207.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/
208.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/209.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/
210.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/211.tif.

Thimbleby:1989:RDC

[Thi89] Harold Thimbleby. A review of Donald C. Lindsay’s text file difference utility, diff. Communications of the Association for Computing Machinery, 32(6):756–755, June 1989. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [VL89].

Thimbleby:2003:ECP

[Thi03] Harold Thimbleby. Explaining code for publication. Software—Practice and Experience, 33(10):975–1001, August 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tung:1989:SMLa

[Tun89a] Sho-Huan Tung. A structured method for literate programming. Technical report 269, Indiana University, Computer Science Department, Bloomington, IN, USA, January 1989. 14 pp.

Tung:1989:SMLb

[Tun89b] Sho-Huan Tung. A structured method for literate programming.

Structured Programming, 10(2):113– 120, 1989. CODEN STPGEM. ISSN 0935-1183.

USENIX:1997:ATT

[USE97] USENIX, editor. 5th Annual Tcl/Tk Workshop ’97, July 14–17, 1997. Boston, MA. USENIX, Berkeley, CA, USA, July 14–17, 1997.

Ammers:1990:LPV

[vA90] E. W. van Ammers. Literate programming with VAMP. Informatie, 32(4):380–388, ???? 1990. CODEN INFTCR. ISSN 0019-9907.

Ammers:1992:VTL

[vAK92a] E. W. van Ammers and M. R. Kramer. VAMP: a tool for literate programming independent of programming language and formatter. In Patrick Dewilde and Joos Vandewalle, editors, CompEuro 1992 Proceedings: Computer Systems and Software Engineering, May 4–8, 1992, The Hague, The Netherlands, pages 371–376. IEEE CS Press, Los Alamitos, CA, USA, May 1992. ISBN 0-8186-2760-3. LCCN QA76.5.C6122 1992.

vanAmmers:1992:VTL

[vAK92b] E. W. van Ammers and M. R. Kramer. VAMP: a tool for literate programming independent of programming language and formatter. In P. Dewilde and J. VandeWalle, editors, CompEuro 1992 Proceedings. Computer Systems and Software engineering, pages 371–376. ????, ????, 1992. ISBN ???? LCCN ????

VandenBosch:1990:WPL

[Van90a] P. N. Van den Bosch. Weaving a program: Literate programming in WEB. ACM Computing Reviews, 31 (7):343–344, 1990.

Wyk:1990:LPA

[Van90b] Christopher J. Van Wyk. Literate programming—an assessment. Communications of the Association for Computing Machinery, 33(3): 361, 365, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanLeeuwen:1995:LPC

[van95] Marc A. van Leeuwen. Literate programming in C: CWEBx manual. Report AM-R9510, Centrum voor Wiskunde en Informatica, Department of Analysis, Algebra and Geometry, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, 1995.

Deursen:1996:LPA

[vDHK96] Arie van Deursen, Jan Heering, and Paul Klint, editors. Language Prototyping: An Algebraic Specification Approach. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2732-9. xx + 353 pp. LCCN QA76.7 .L355 1996. URL http:

//www.cwi.nl/~gipe/language_
prototyping/index.html. Many of

the chapters in this book are literate programs.

Vee:2017:CLH

[Vee17] Annette Vee. Coding Literacy: How Computer Programming is Changing Writing. Software studies. MIT Press, Cambridge, MA, USA, 2017. ISBN 0-262-03624-X (hardcover). xi + 361 pp. LCCN LC149.5.

VanWyk:1988:LPE

[VHC88] Christopher J. Van Wyk, Eric Hamilton, and Don Colner. Literate programming: Expanding generalized regular expressions. Communications of the Association for Computing Machinery, 31(12):1376– 1385, December 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

VanWyk:1987:LPPa

[VHG87] Christopher J. Van Wyk, David R. Hanson, and John Gilbert. Literate programming: Printing common words. Communications of the Association for Computing Machinery, 30(7):594–599, July 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL ftp://ftp.cs.princeton.
edu/pub/packages/loom.tar.Z. See

another solution to the problem of printing the k most common words of a text file in [BKM86].

VanWyk:1987:LPPb

[VJW87] Christopher J. Van Wyk, Michael Jackson, and David W. Wall. Literate programming: Processing transactions. Communications of the Association for Computing Machinery, 30(12):1000–1010, December 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

VanWyk:1989:LPF

[VL89] Christopher J. Van Wyk and Donald C. Lindsay. Literate programming: a file difference program. Communications of the Association for Computing Machinery, 32(6):740–755, June 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See review [Thi89].

Vestdam:2002:AIP

[VN02] Thomas Vestdam and Kurt Nørmark.

Aspects of internal program documentation: An elucidative perspective. In IEEE, editor, Proceedings of the 10th International Workshop on Program Comprehension, June 27–29, 2002, Paris, France, pages 43–52. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1495-2 (paperback), 0-7695-1496-0 (case), 0-7695-1497-9 (microfiche). ISSN 1092-8138. LCCN

???? URL http://www.gbv.de/
dms/bowker/toc/9780769514963.

IEEE Computer Society order number PR01495.

Vestdam:2005:TDP

[VN05] T. Vestdam and K. Nørmark. Toward documentation of program evolution. In IEEE, editor, Proceedings of the 21st IEEE International Conference on Software Maintenance: [Budapest, Hungary, 26 - 29 September 2005]. ICSM’05. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005.

ISBN 0-7695-2368-4. ISSN 1063-6773. LCCN QA76.76.S64.

VanWyk:1989:LPW

[VR89] Christopher J. Van Wyk and Norman Ramsey. Literate programming: Weaving a languageindependent WEB. Communications of the Association for Computing Machinery, 32(9):1051– 1055, September 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waite:1993:ELD

[Wai93] W. M. Waite. An executable language definition. ACM SIG-PLAN Notices, 28(2):21–40, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper describes a subset of C, called C–, defined as a literate program using the Eli and FunnelWeb systems.

Wu:1989:SCD

[WB89] Y. C. Wu and Ted P. Baker. A source code documentation system for Ada. ACM Ada Letters, 9(5):84– 88, 1989.

Williams:1992:FUM

[Wil92] Ross Williams. FunnelWeb user’s

                                        manual.  ftp.adelaide.edu.au
                                        in /pub/compression and /pub/

funnelweb, University of Adelaide, Adelaide, South Australia, Australia, 1992.

Williams:2000:FTM

[Wil00] R. Williams. FunnelWeb tutorial manual. Web document,

2000.  URL http://www.ross.
net/funnelweb/tutorial/intro_
what.html.

Wittenberg:1995:LPC

[Wit95] Lee Wittenberg. Literate programming in C and C++ using CWEB. C/C++ Users Journal, 13(5):55, 57–63, May 1995. CODEN CCU-JEX. ISSN 1075-2838.

Wittenberg:1996:ULPa

[Wit96a] L. Wittenberg. Using literate programming notation in introductory programming courses. In Rosenberg [Ros96], pages 267–272. ISBN 0-89791-845-2. LCCN ???? ACM Order no. 457963.

Wittenberg:1996:ULPb

[Wit96b] L. Wittenberg. Using literate programming notation in introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):437, ???? 1996. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wittenberg:1996:ULP

[Wit96c] Lee Wittenberg. Using literate programming notation in introductory programming courses. In Proceedings of the 1st Australasian conference on computer science education, Sydney, Australia, pages 267–272. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-845-2. LCCN ????

Wu:1990:WLP

[Wu90] Cheng-Shiung Jesse Wu. A wysiwyg literate programming system. Thesis (Ph.D.), Dept. of Computer and Information Science, Ohio State University, Columbus, OH, USA, 1990. vii + 51 pp. Thesis advisor: Eitan Gurari.

Zeng:1991:LPS

[Zen91] Yunfeng Zeng. Literate programming system CDS. Journal of Computer Science and Technology, 6 (3):263–270, July 1991. CODEN JCTEEM. ISSN 1000-9000.

Zukowski:1997:ISU

[Zuk97] Monty Zukowski. Implementing a selective undo framework in Python. In Anonymous, editor, Proceedings of the 6th International Python Conference: October 14–17, 1997, San Jose, California, pages 69– 75. Corporation for National Research Initiatives (CNRI), 1895 Preston White Drive, Suite 100, Reston, VA 20191, USA, 1997.

                                        LCCN ???? URL http://www.
                                        python.org/workshops/1997-10/
                                        proceedings/zukowski.html.